### 15 Mar 2026 v1

# A Unified Pulse-Shaped OFDM Framework for Chirp-Domain Waveforms: Continuous-Time Modeling and Practical I/O Analysis

Abstract
In this paper, a unified framework for chirp-domain waveforms, including orthogonal chirp division multiplexing (OCDM) and affine frequency division multiplexing (AFDM), is developed. Based on their continuous-time representations, we show that these waveforms fall within the conventional Weyl-Heisenberg (WH) framework for multicarrier (MC) waveforms, where the root chirp corresponds directly to the prototype pulse in the WH framework. Since the chirp is a constant-envelope signal and is transparent to subcarrier orthogonality, these waveforms can be further interpreted as pulse-shaped (PS) orthogonal frequency division multiplexing (OFDM). Within the developed PS-OFDM framework, the power spectral density of chirp-domain waveforms is derived analytically. We then discuss existing practical implementations of chirp-domain waveforms, which rely on sub-Nyquist discrete-time samples and therefore exhibit frequency aliasing. The resulting aliased waveform is analyzed, and the orthogonality among the embedded aliased chirps is discussed. It is shown that the aliased chirps are conditionally orthogonal, whereas the implemented approximate aliased chirps can maintain mutual orthogonality when an appropriate sample-wise pulse-shaping filter is applied. We further derive an exact input-output relation for the implemented chirp-domain waveform over a delay-Doppler (DD) channel, showing that the effective channel observed at a practical receiver does not, in general, admit a DD spreading-function model commonly assumed in the literature. The implementation complexity is also investigated and compared with that of orthogonal delay-Doppler division multiplexing (ODDM), the DD-domain MC waveform defined within the evolved WH framework. Finally, simulation results are provided to verify the analysis.
I Introduction
Next-generation wireless systems, including sixth-generation (6G) mobile networks, vehicle-to-everything (V2X) communications, high-speed railway links, and millimeter-wave/terahertz (mmWave/THz) systems, are increasingly expected to operate in scenarios with very high user or scatterer mobility. Such mobility induces a large Doppler spread and shortens the channel coherence time. In conventional orthogonal frequency-division multiplexing (OFDM) systems, this often manifests as rapidly increasing inter-carrier interference (ICI): once the normalized Doppler spread (e.g., the maximum Doppler shift divided by the subcarrier spacing) is no longer sufficiently small, subcarriers may effectively lose their frequency-domain orthogonality. This behavior is consistent with the fact that OFDM is typically designed under an approximately linear time-invariant (LTI) channel assumption, whereas under high mobility the physical channel is more appropriately modeled as a linear time-varying (LTV) system. Consequently, it becomes necessary to represent the channel in domains that better capture the LTV structure, and to design waveforms and transceivers accordingly.
Within the classical wide-sense stationary uncorrelated scattering (WSSUS) model[1, 2, 3], an LTV channel is statistically characterized by the scattering function, whose marginals yield the power-delay profile (PDP) and the Doppler power spectrum. However, in highly time-varying wideband regimes, these statistical averages may not explicitly reveal how transmitted symbols are influenced by individual propagation paths. To obtain a more structured view, it is useful to adopt a delay-Doppler (DD) representation, in which the channel is described by a DD spreading function that quantifies how signal energy is shifted by each path’s delay and Doppler before reaching the receiver. This representation makes channel effects more transparent, thereby motivating waveform and equalization designs that explicitly account for the channel structure in the DD domain.
Recent waveform designs for LTV channels can be broadly categorized, from the DD perspective, into two representative approaches. The first is straightforward DD-domain waveforms, which aim to directly exploit structured DD-domain channel behavior and the associated diversity by multiplexing symbols on a DD grid, as exemplified by orthogonal time-frequency space (OTFS) modulation [4] and orthogonal delay-Doppler multiplexing (ODDM) [5]. The second approach consists of chirp-domain waveforms, including orthogonal chirp division multiplexing (OCDM) [6] and its generalization affine frequency division multiplexing (AFDM) [7], which adopt a chirp-domain view by employing chirp-based subcarriers instead of sinusoidal ones in OFDM. Since chirps exhibit a swept instantaneous frequency, a path-induced delay shift can be translated into a frequency shift and combined with the Doppler effect of the path. Such chirp-domain waveforms aim to improve robustness to the channel’s DD effects while maintaining an OFDM-like block structure and comparable implementation complexity.
Since the physical units of delay and Doppler are time and frequency, respectively, the DD domain is naturally a time-frequency (TF) domain. As a result, it has been shown that DD-domain waveforms can be interpreted within the Weyl-Heisenberg (WH) framework for multicarrier (MC) waveforms, where ODDM evolves the existing WH framework by introducing a (bi)orthogonal WH-subset-based design [8, 9]. Recent studies of ODDM waveforms over general physical channels can also be found in [10, 11, 12].
On the other hand, although orthogonal chirps play a central role in chirp-based waveforms, a systematic framework for understanding the fundamentals of these orthogonal-chirp-based waveforms is still lacking. Moreover, existing studies of chirp-domain waveforms are often based on discrete-time sequences generated by the discrete Fresnel transform (DFnT) [6] or the discrete affine Fourier transform (DAFT)[7], whereas the transmitted waveform is ultimately a continuous-time signal that propagates through a general physical channel with continuous-valued delay and Doppler. This discrepancy may lead to a mismatch between discrete-sequence-level analyses and the underlying channel behavior characterized by continuous-valued delay and Doppler.
In this paper, we develop a unified framework for chirp-domain waveforms, including OCDM and AFDM. We show that these waveforms also fall within the conventional WH framework for MC waveforms, where the root chirp corresponds directly to the prototype pulse in the WH framework. Under this interpretation, chirp-domain waveforms can be viewed as pulse-shaped (PS) OFDM with a constant-envelope prototype pulse. We also analyze practical implementations based on sub-Nyquist discrete-time samples, investigate the resulting aliased chirps and their orthogonality properties, and derive the exact input-output (I/O) relation over DD channels. The contributions of this paper are summarized as follows:
  • Based on their continuous-time representations, we show that chirp-domain waveforms belong to the conventional WH framework for MC waveforms. In this formulation, the root chirp corresponds to the prototype pulse, and chirp-domain waveforms can be interpreted as PS-OFDM signals with a constant-envelope prototype pulse, which is transparent to subcarrier orthogonality. The developed unified PS-OFDM framework connects chirp-domain waveform design with classical MC signal theory and provides a systematic perspective for analyzing their power spectral density (PSD).
  • Building on the derived PSD, we investigate practical implementations of chirp-domain waveforms based on sub-Nyquist discrete-time samples. This leads to the concept of aliased chirps, whose orthogonality properties are analyzed. We show that the aliased chirps are conditionally orthogonal, whereas the approximate aliased chirps used in practical implementations can maintain mutual orthogonality when appropriate sample-wise pulse-shaping filters are applied.
  • For implemented waveforms containing approximate aliased chirps, we derive the exact I/O relation over DD channels. Our analysis shows that the effective channel observed at the receiver does not, in general, admit the conventional DD spreading-function representation commonly assumed in the literature. As a result, compared with existing studies that rely solely on discrete-sequence models and may therefore deviate from the behavior observed in practical transceivers, the derived I/O relation provides a more accurate characterization of the practical performance of chirp-domain waveforms.
  • Simulations are provided to verify the theoretical analysis, validate the derived PSD of chirp-domain waveforms, and confirm the derived I/O relation. We also analyze the implementation complexity of chirp-domain waveforms and compare it with that of ODDM. When occupying the same TF resources, the implementation complexity of ODDM is shown to be significantly lower than that of chirp-domain waveforms.
The remainder of this paper is organized as follows. In Section II, beginning with a general representation of ditigal modulation waveforms, the WH framework for MC waveforms is introduced. Section III presents OCDM and AFDM in their continuous-time forms. Based on this, the unified PS-OFDM framework for chirp-domain waveforms is developed and the PSD is derived. In Section IV, practical implementations based on sub-Nyquist samples are investigated; the ideal and approximate aliased chirps are presented, and their orthogonality properties are studied. Section V derives the I/O relation of implemented chirp-domain waveforms with approximate aliased chirps over DD channels and verifies the effective channel observed at the receiver. Simulation validation and complexity comparison are also provided in Section VI. Finally, Section VII concludes the paper.
Notations: In this paper, uppercase boldface letters are used to represent matrices, and lowercase boldface letters are used for column vectors. The superscript $\mathsf{T}$, $\mathsf{H}$ and * denote the transpose, conjugate transpose, and conjugate operators, respectively. Also, $\Pi_{T}(t)$ denotes the rectangular function of unit energy and support of $[0,T)$, while $\mathcal{A}_{g,\tilde{g}}(\tau,\nu)$ stands for the cross ambiguity function between two pulses $g(t)$ and $\tilde{g}(t)$, given by
$\mathcal{A}_{g,{\tilde{g}}}(\tau,\nu)$ $=\langle g(t),{\tilde{g}}(t-\tau)e^{j2\pi\nu(t-\tau)}\rangle,$
$=\int_{-\infty}^{+\infty}g(t){\tilde{g}}^{*}(t-\tau)e^{-j2\pi\nu(t-\tau)}dt.$
II MC Waveform Principles
Given the symbol index $i\in\mathbb{Z}$, a real-valued passband digital modulation waveform with carrier frequency $f_{c}$ and phase $\theta_{c}$ is typically represented as [13, 14, 15, 9]:
$x_{pb}(t)=\sum_{i}\sqrt{2}A_{i}\cos(2\pi(f_{c}+f_{i}(t))t+\theta_{c}+\theta_{i})\rho_{i}(t),$
where the amplitude $A_{i}$, the phase $\theta_{i}$, the possibly time-varying frequency $f_{i}(t)$, and the finite-duration pulse $\rho_{i}(t)$ can be used for modulation. The corresponding complex-valued baseband waveform is $x(t)=\sum_{i}x_{i}(t)$, where the $i$-th symbol is defined as:
$x_{i}(t)=A_{i}e^{j\theta_{i}}e^{j2\pi f_{i}(t)t}\rho_{i}(t)\triangleq X[i]g_{i}(t).$ (1)
This consists of two components: an information-bearing digital symbol (a complex number) $X[i]=A_{i}e^{j\theta_{i}}$, drawn from a signaling alphabet, and a finite-energy, continuous-time function $g_{i}(t)=e^{j2\pi f_{i}(t)t}\rho_{i}(t)$, referred to as the transmit (modulating) pulse or filter.
In the receiver, a correlator performing cross-correlation with a receive pulse $\gamma_{i}(t)$ is typically used to extract the $i$-th digital symbol as:
$Y[i]=\int_{-\infty}^{+\infty}y(t)\gamma_{i}^{*}(t)dt,$ (2)
where $y(t)$ is the received waveform. The correlator can be implemented by a matched filter [13, 14, 15], whose output sample at an appropriate time yields $Y[i]$.
To avoid inter-symbol interference (ISI), the transmit pulses $\{g_{i}(t),i\in\mathbb{Z}\}$ are required to be mutually orthogonal. Hence, a digital modulation waveform is defined by a set of orthogonal pulses, also referred to as basis functions.
For MC systems, the symbol index $i$ is replaced by a time index $m$ and a frequency index $n$, and the MC waveform is given by [8, 9, 16]:
$x(t)=\sum_{m=0}^{M-1}\sum_{n=-\frac{N}{2}}^{\frac{N}{2}-1}X[m,n]\overbrace{e^{j2\pi n\Delta F(t-m\Delta T)}g(t-m\Delta T)}^{g_{m,n}(t)},$ (3)
where $N$ and $M$ are the numbers of subcarriers and MC symbols, respectively. $\Delta T$ and $\Delta F$ are the time and frequency resolutions, and $g(t)$ is called the prototype pulse. In the literature, $\Delta T$ is also known as the symbol interval, while $\Delta F$ is referred to as the fundamental frequency or subcarrier spacing. Since $e^{j2\pi n\Delta Ft}$ is often called a subcarrier or tone, the transmit pulses $g_{m,n}(t)$ in (3) are truncated (windowed) or pulse-shaped subcarriers, corresponding to a rectangular or non-rectangular $g(t)$, respectively.
For such an MC waveform characterized by a TF grid $(\Delta T,\Delta F)$ and a prototype pulse $g(t)$, a design framework based on the WH frame theory is well-known[16], as $g_{m,n}(t)$ in (3) is the WH or Gabor function adopted in TF signal analysis [17].
A WH set $\{\Delta T,\Delta F,g(t)\}$ is said to be orthogonal given the inner product:
$\langle g_{m,n}(t),g_{m^{\prime},n^{\prime}}(t)\rangle=\delta(m-m^{\prime})\delta(n-n^{\prime}).$ (4)
Let the corresponding receive pulse be another WH function $\gamma_{m,n}(t)=\gamma(t-m\Delta T)e^{j2\pi n\Delta F(t-m\Delta T)}$. A pair of WH sets $\{\Delta T,\Delta F,g(t)\}$ and $\{\Delta T,\Delta F,\gamma(t)\}$ is said to be biorthogonal when:
$\langle g_{m,n}(t),\gamma_{m^{\prime},n^{\prime}}(t)\rangle=\delta(m-m^{\prime})\delta(n-n^{\prime}).$ (5)
According to the WH frame theory, (bi)orthogonal WH sets only exist when the joint TF resolution $\Delta R=\Delta T\Delta F\geq 1$. A typical example of such a (bi)orthogonal WH set is that adopted in OFDM, with $\Delta T=T+T_{cp}$, $\Delta F=1/T$, $g(t)=\Pi_{T+T_{cp}}$, and $r(t)=\Pi_{T}$, where $T_{cp}$ is the length of the cyclic prefix (CP). MC waveforms employing (bi)orthogonal WH functions with non-rectangular $g(t)$ are therefore called PS-OFDM[18].
III PS-OFDM Framework for OCDM/AFDM
III-A OCDM Waveform
An OCDM signal having $N$ orthogonal chirps can be written as [6]
$x_{\mathrm{OC}}(t)=\sum_{n=0}^{N-1}X[n]\,\psi_{n}(t),\quad 0\leq t<T,$ (6)
where $X[n]$ is the information-bearing complex symbol,
$\psi_{n}(t)=\Pi_{T}(t)e^{j\frac{\pi}{4}}e^{-j\alpha\pi\left(t-\frac{n}{\alpha T}\right)^{2}},$ (7)
is the $n$th chirp and $\alpha=\frac{N}{T^{2}}$ is the chirp rate. The corresponding root chirp is given by
$\psi_{0}(t)=\Pi_{T}(t)e^{j\frac{\pi}{4}}e^{-j\alpha\pi t^{2}}.$ (8)
It is known that these chirps are orthogonal to each other, evidenced by
$\int_{0}^{T}\psi_{n}(t)\,\psi^{*}_{n^{\prime}}(t)\,dt=T\delta(n-n^{\prime}).$ (9)
Without loss of generality, let NN be an even number. The implementation of OCDM signal is based on the DFnT[6], as the discrete OCDM signal is given by sampling the waveform in (6) at a rate of $\frac{N}{T}$ as
$x_{\rm{OC}}[k]$ $=x_{\rm{OC}}\left(k\frac{T}{N}\right)$
$=\sum_{n=0}^{N-1}X[n]e^{\frac{j\pi}{4}}e^{-j\frac{\pi}{N}(k-n)^{2}},$ (10)
for $0\leq k\leq N-1$, where $e^{\frac{j\pi}{4}}e^{-j\frac{\pi}{N}(k-n)^{2}}$ is exactly the $(k,n)$-th entry of the $N\times N$ IDFnT matrix up to a scaling factor.
III-B AFDM Waveform
The foundation of the AFDM signal is the affine Fourier transform (AFT) namely linear canonical transform (LCT) [19], given by
$X(u)=\int_{-\infty}^{+\infty}x(t)\mathcal{K}(t,u)dt,$ (11)
where the kernel
$\mathcal{K}(t,u)=\frac{1}{\sqrt{2\pi|b|}}e^{-j\left(\frac{a}{2b}u^{2}+\frac{1}{b}ut+\frac{d}{2b}t^{2}\right)},$ (12)
has four parameters $(a,b,c,d)$ that satisfy the constraint adbc=1ad-bc=1. Without loss of generality, we assume $b>0$ hereafter. It is known that $x(t)$ can be recovered as
$x(t)=\int_{-\infty}^{+\infty}X(u)\mathcal{K}^{*}(t,u)du.$ (13)
using the inverse AFT, which is also an AFT with parameters (a,b,c,d)(a,-b,-c,d).
By sampling (11) over both $t$ and $u$ with intervals $\Delta t$ and $\Delta u$, respectively, the discrete AFT (DAFT) and the corresponding inverse DAFT can be defined under the condition [20]
$\Delta t\Delta u=\frac{2\pi b}{N},$ (14)
where $N$ denots denotes the number of samples over $u$. For the commonly adopted symmetric sampling grid with $\Delta t=\Delta u$, the transform parameter $b$ follows
$b=\frac{N\Delta u^{2}}{2\pi}.$ (15)
In [7], the inverse DAFT is used to generate the discrete-time AFDM sequence as
$x_{\rm{AF}}[k]$ $=\frac{1}{\sqrt{N}}\sum_{n=0}^{N-1}X[n]e^{j2\pi\left(c_{2}n^{2}+\frac{nk}{N}+c_{1}k^{2}\right)}$ (16)
for $0\leq k\leq K-1$, where $K=N$, $X[n]\triangleq X(n\Delta u)$ and
$c_{1}$ $=\frac{d}{4\pi b}\Delta t^{2},$ (17)
$c_{2}$ $=\frac{a}{4\pi b}\Delta u^{2}.$ (18)
Comparing (16) with (13), one can see that the ideal continuous-time AFDM signal can be obtained by considering sampling over $u$ only. Substituting $u=n\Delta u$ into (13) for $0\leq t<T$ and replacing the integral by a Riemann sum yield
$x_{\rm{AF}}(t)\approx\sum_{n=0}^{N-1}X[n]\underbrace{\mathcal{K}^{*}(t,n\Delta u)\Delta u}_{\triangleq\dot{\phi}_{n}(t)},$ (19)
where
$\dot{\phi}_{n}(t)$ $=\frac{\Delta u}{\sqrt{2\pi|b|}}\Pi_{T}(t)e^{j\left(\frac{a}{2b}(n\Delta u)^{2}+\frac{1}{b}(n\Delta u)t+\frac{d}{2b}t^{2}\right)},$
$=\frac{1}{\sqrt{N}}\Pi_{T}(t)e^{j2\pi\left(c_{2}n^{2}+\frac{n}{N\Delta t}t+c_{1}\left(\frac{t}{\Delta t}\right)^{2}\right).}$ (20)
Bearing in mind that $\Delta t=\frac{T}{N}$, and comparing (20) with (7), it is clear that OCDM is a special case of AFDM with $c_{1}=\frac{1}{2N}.$ Thus, we hereafter use
$x_{\rm{AF}}(t)=\sum_{n=0}^{N-1}X[n]\dot{\phi}_{n}(t)$ (21)
as the general form of chirp-domain waveforms for subsequent discussion.
III-C WH and PS-OFDM Frameworks
By absorbing the term $\frac{1}{\sqrt{N}}e^{j2\pi c_{2}n^{2}}$ into $X[n]$ as $\dot{X}[n]=\frac{1}{\sqrt{N}}e^{j2\pi c_{2}n^{2}}X[n]$, (21) can be rewritten as
$x_{\rm{AF}}(t)=\sum_{n=0}^{N-1}\dot{X}[n]\phi_{n}(t),$ (22)
where
$\phi_{n}(t)$ $=\Pi_{T}(t)e^{j2\pi c_{1}N^{2}\left(\frac{t}{T}\right)^{2}}e^{j2\pi\frac{n}{T}t}.$ (23)
Comparing (22) with (3), one can see that (22) corresponds exactly to an MC symbol, and the basis functions in (23) take the form of WH functions, where $\Delta T=T$, $\Delta F=1/T$, and the prototype pulse becomes
$g_{\rm AF}(t)$ $=\Pi_{T}(t)e^{j2\pi c_{1}N^{2}\left(\frac{t}{T}\right)^{2}},$ (24)
which is the root chirp up to a scaling factor.
It is interesting to observe that this chirp-based prototype pulse is a constant-envelope signal, which does not affect the orthogonality among the basis functions, because
$\int_{0}^{T}\phi_{n}(t)\phi^{*}_{n^{\prime}}(t)\,dt$ $=\int_{0}^{T}g_{\rm AF}(t)g^{*}_{\rm AF}(t)e^{j2\pi\frac{n-n^{\prime}}{T}t}dt,$
$=\int_{0}^{T}e^{j2\pi\frac{n-n^{\prime}}{T}t}dt,$
$=T\delta(n-n^{\prime}).$ (25)
In other words, $g_{\rm AF}(t)$ can form an orthogonal WH set over the TF grid $\left(T,\frac{1}{T}\right)$.
Recall that when $g(t)$ forms an orthogonal WH set over the TF grid $\left(\Delta T,\Delta F\right)$, the MC waveform in (3) becomes a PS-OFDM signal. As a result, chirp-domain waveforms not only belong to the WH framework but can also be interpreted as PS-OFDM signals with a constant-envelope prototype pulse.
III-D PSD Analysis
Within the above unified PS-OFDM framework, the PSD of chirp-domain waveforms can be derived analytically.
III-D1 PSD of PS-OFDM Signals
Assume that the information-bearing symbols $X[m,n]$ are i.i.d. with zero mean and average power $\sigma_{X}^{2}=\mathbb{E}[|X[m,n]|^{2}]$. It can be proved that $x(t)$ is wide-sense cyclostationary with period $\Delta T$. Given the time-average autocorrelation function of cyclostationary signal is defined as [14] :
$\bar{R}_{x}(\tau)$ $=\frac{1}{\Delta T}\int_{0}^{\Delta T}R_{x}(t,\tau)dt,$
$=\frac{1}{\Delta T}\int_{0}^{\Delta T}\mathbb{E}[x(t)x^{*}(t-\tau)]dt,$ (26)
using (3), we have
$\mathbb{E}[x(t)x^{*}(t-\tau)]=$ $\sigma_{X}^{2}\sum_{m}g(t-m\Delta T)g^{*}(t-\tau-m\Delta T)$
$\times\sum_{n}e^{j2\pi n\Delta F\tau}.$ (27)
Substituting (27) into (III-D1), we obtain
$\bar{R}_{x}(\tau)$ $=\frac{\sigma_{X}^{2}}{\Delta T}R_{g}(\tau)\sum_{n=-\frac{N}{2}}^{\frac{N}{2}-1}e^{j2\pi n\Delta F\tau},$ (28)
where $R_{g}(\tau)$ is the autocorrelation of the prototype pulse given by
$R_{g}(\tau)=\int_{-\infty}^{+\infty}g(t)\,g^{*}(t-\tau)\,dt.$ (29)
Applying the Wiener-Khinchin theorem, we have the PSD
$S_{x}(f)$ $=\mathcal{F}\{\bar{R}_{x}(\tau)\},$
$=\frac{\sigma_{X}^{2}}{\Delta T}\sum_{n=-\frac{N}{2}}^{\frac{N}{2}-1}|G(f-n\Delta F)|^{2},$ (30)
where $G(f)$ is the Fourier transform of $g(t)$. (30) shows that the PSD is determined by the prototype pulse spectrum $|G(f)|^{2}$, the frequency resolution $\Delta F$, and the time resolution $\Delta T$.
III-D2 PSD of AFDM Waveform
For the ideal AFDM signal in (22), after a centered re-indexing of the subcarrier index to match the unified formulation in (3), and then substituting $g(t)=g_{\rm AF}(t)$, $\Delta T=T$, $\Delta F=1/T$ into (30), we have
$S_{\rm AF}(f)=\frac{\sigma_{X}^{2}}{NT}\sum_{n=-\frac{N}{2}}^{\frac{N}{2}-1}\left|G_{\rm AF}\left(f-\frac{n}{T}\right)\right|^{2},$ (31)
where $e^{j2\pi c_{2}n^{2}}$ is a phase factor that does not affect the second-order statistics of $\dot{X}[n]$, and $G(f)$ is the corresponding prototype pulse spectrum given by
$G_{\rm AF}(f)=\int_{0}^{T}e^{j2\pi c_{1}N^{2}\left(\frac{t}{T}\right)^{2}}e^{-j2\pi ft}\,dt.$ (32)
Let the span of $G_{\rm AF}(f)$ be $B_{g_{\rm AF}}$, which can be considered as the bandwidth of $g_{\rm AF}(t)$ [^1] The bandwidth can be reasonably defined by ignoring the negligibly small high-frequency tails beyond $B_{g_{\rm AF}}/2,B_{g_{\rm AF}}/2]$.. Without loss of generality, we can frequency-shift the root chirp by $e^{-j2\pi\frac{c_{1}N^{2}}{T}t}$. Then, the instantaneous frequency of the root chirp ranges from $\frac{c_{1}N^{2}}{T}$ to $\frac{c_{1}N^{2}}{T}$, and we obtain $B_{g_{\rm AF}}\approx\frac{2c_{1}N^{2}}{T}$ by ignoring the edge effect of $\Pi_{T}(t)$. From (32), the bandwidth of the ideal AFDM signal is given by
$B_{\rm AF}=\frac{N-1}{T}+B_{g_{\rm AF}}\approx\frac{2c_{1}N^{2}+N-1}{T}.$ (33)
IV Aliased AFDM Waveforms
Recall that the AFDM sequence has sampling interval $\Delta t=\frac{T}{N}$, or equivalently a sampling frequency of $\frac{N}{T}$. Because $B_{\rm AF}>\frac{N}{T}$, the AFDM sequence generated by the DAFT consists of sub-Nyquist samples of the ideal AFDM waveform, which results in frequency aliasing.
IV-A Ideal Aliased AFDM Waveform
Let $C=2N|c_{1}|$, the ideal aliased AFDM waveform is composed of aliased chirps [21, 22]
$\hat{\phi}_{n}(t)=e^{j2\pi\left(c_{2}n^{2}+\frac{c_{1}}{\Delta t^{2}}t^{2}+\frac{n}{T}t-\frac{q}{\Delta t}t\right)},$ (34)
for $t_{n,q}\leq t<t_{n,q+1}$, with
$t_{n,q}=\begin{cases}0,&q=0,\\[4.0pt] \frac{T}{C}q-\frac{n}{C}\Delta t,&q=1,2,\ldots,C,\\[6.0pt] T,&q=C+1.\end{cases}$ (35)
However, these aliased chirps are conditionally orthogonal (see Appendix A for proof). Consequently, interference may arise in the aliased AFDM waveform, rendering it difficult for communication applications.
IV-B Approximate Aliased AFDM Waveform
On the other hand, let $a(t)$ denote the pulse-shaping filter. The aliased AFDM waveform may be approximated by applying an $a(t)$-based sample-wise pulse shaping to the AFDM sequence in (16), or equivalently by passing (16) through $a(t)$.
Remark 1
A discrete sequence has infinite bandwidth, and therefore must be pulse-shaped into a analog waveform that is band-limited before being carrier-modulated and transmitted over the air.
Remark 2
In the context of MC waveforms, pulse shaping can be somewhat confusing because an MC symbol usually spans multiple time-domain samples, whereas in single-carrier modulation one baud-rate sample corresponds to one symbol. Pulse shaping in MC waveforms typically involves two steps: first, interpolating discrete-time samples using a filter (equivalent to the sample-wise pulse shaping used in the single-carrier case), and then multiplying the result by the symbol-wise prototype pulse (symbol-wise pulse shaping). When the prototype pulse is constant-envelope, such as the rectangular pulse, sample-wise pulse shaping alone may be sufficient, subject to appropriate parameter settings [8, 9].
After the sample-wise pulse shaping, we have the approximate aliased AFDM waveform
$x_{\rm AF}^{(a)}(t)=\sum_{k}x_{\rm AF}[k]a\left(t-k\frac{T}{N}\right),$ (36)
and the corresponding approximate aliased chirps are given by
$\phi_{n}^{(a)}(t)$ $=\sum_{k=0}^{N-1}\phi_{n}[k]\,a\left(t-k\frac{T}{N}\right),$ (37)
where the $n$-th chirp sequence is
$\phi_{n}[k]$ $=\phi_{n}\!\left(k\frac{T}{N}\right)=e^{j2\pi c_{1}k^{2}}e^{j2\pi\frac{nk}{N}}$ (38)
for $0\leq k\leq N-1$. The chirp sequence can be represented as a vector $\bm{\phi}_{n}=[\phi_{n}[0],\cdots,\phi_{n}[N-1]]^{T},$ corresponding to the $n$-th column of the IDAFT matrix. Since $\langle\bm{\phi}_{n},\bm{\phi}_{n^{\prime}}\rangle=\delta(n-n^{\prime})$, it is well known that
$\langle\phi_{n}^{(a)}(t),\phi_{n^{\prime}}^{(a)}(t)\rangle=\delta(n-n^{\prime}),$
when $a(t)$ is a root-Nyquist pulse for the Nyquist interval $\frac{T}{N}$ [15].
As a result, to maintain orthogonality among these approximated aliased chirps, a root-Nyquist pulse-shaping filter is required.
V I/O Relation over DD channels
Assume the LTV channel consists of $P$ propagation paths. For the $p$-th path, let $h_{p}$, $\tau_{p}$, and $\nu_{p}$ denote its complex gain, delay, and Doppler shift, respectively. This channel model is considered valid during the channel’s ”stationarity” interval[1, 3, 8]. Its representation in the DD domain, also known as the DD spreading function, is given by
$\tilde{h}(\tau,\nu)=\sum_{p=1}^{P}h_{p}\delta(\tau-\tau_{p})\delta(\nu-\nu_{p}),$ (39)
where $\tau$ and $\nu$ are the delay and Doppler domain variables, respectively. Without loss of generality, we assume
$\tau_{1}\leq\tau_{2}\leq\cdots\leq\tau_{P}.$ (40)
and the delay spread of the channel is given by $\mathcal{T}=\tau_{p}-\tau_{1}<T$. For a transmitted baseband signal $x(t)$, the received baseband signal $r(t)$ is expressed as
$y(t)=\sum_{p=1}^{P}h_{p}e^{j2\pi\nu_{p}(t-\tau_{p})}x(t-\tau_{p})+z(t),$ (41)
where $z(t)$ represents additive white Gaussian noise (AWGN) with a PSD of $N_{0}$.
V-A Discrete-Sequence Based I/O Relation
The I/O relation of AFDM waveform has been derived in [7], where each path is assumed with an on-grid delay that can be normalized to an integer as $l_{p}=\tau_{p}\frac{N}{T}.$ Then, the received samples are given as [7]
$y[k^{\prime}]$ $=\sum_{l=0}^{\infty}h_{k^{\prime},l}\,x_{\rm AF}[k^{\prime}-l]+z[k^{\prime}],$ (42)
where $z[k^{\prime}]$ is the samples of AWGN and
$h_{k^{\prime},l}$ $=\sum_{p=1}^{P}h_{p}e^{j2\pi\nu_{p}k^{\prime}}\delta(l-l_{p})$ (43)
is the time-varying impulse response of channel at time $k^{\prime}$ and delay $l$. A chirp-periodic prefix (CPP) with length $L_{\rm CPP}>\lceil{\mathcal{T}}\frac{N}{T}\rceil-1$ is ususally prepended into $x_{AF}[k]$ in (42).
After discarding the CPP, (42) can be written in the matrix form
$\mathbf{y}$ $=\mathbf{H}\mathbf{x}+\mathbf{z},$ (44)
with $\mathbf{z}\sim\mathcal{CN}(0,N_{0}\mathbf{I})$,
$\mathbf{H}=\sum_{p=1}^{P}h_{p}\mathbf{\Gamma}_{p}\mathbf{\Delta}_{\nu_{p}}\mathbf{\Pi}^{l_{p}},$ (45)
and $\mathbf{\Pi}$ is the $N\times N$ forward cyclic-shift matrix, $\mathbf{\Delta}_{\nu_{p}}\triangleq\mathrm{diag}\left(1,e^{j2\pi\nu_{p}},\cdots,e^{j2\pi(N-1)\nu_{p}}\right)$ and $\mathbf{\Gamma}_{p}$ is a $N\times N$ diagonal matrix given by
$\mathbf{\Gamma}_{p}$ $=\mathrm{diag}\!\left(\begin{cases}e^{-j2\pi c_{1}\left(N^{2}-2N(l_{p}-n)\right)},&n<l_{p}\\ 1,&n\geq l_{p}\end{cases}\right).$ (46)
By applying the DAFT matrix $\mathbf{A}$ to $\mathbf{y}$, the DAFT domain I/O relation becomes
$\mathbf{y}_{u}$ $=\mathbf{A}\mathbf{y}=\underbrace{\mathbf{A}\mathbf{H}\mathbf{A}^{\mathsf{H}}}_{\mathbf{H}_{u}}\mathbf{x}_{u}+\mathbf{A}\mathbf{z}.$ (47)
where $\mathbf{x}_{u}=[X[0],\cdots,X[N-1]]^{\mathsf{T}}$.
The DAFT domain I/O matrix $\mathbf{H}_{u}$ plays a fundamental role in performance evaluation and AFDM waveform design, such as in selecting the parameters $c_{1}$ and $c_{2}$ to achieve optimal diversity order and reduced pilot overhead [7].
V-B Practical Waveform based I/O Relation
Although the I/O relation in (47) appears appealing, it is derived from the time-domain relation in (42), which, however, does not match the behavior of a practical receiver processing chain, as will be explained below.
The passband counterpart of (41) is given by
$y_{pb}(t)=$ $\sum_{p=1}^{P}\tilde{h}_{p}\cos(2\pi(f_{c}+\nu_{p})(t-\tau_{p}))x(t-\tau_{p})$
$+y_{pb}^{I}(t)+z_{pb}(t),$ (48)
where $\tilde{h}_{p}$ is a real-valued gain and $h_{p}=\tilde{h}_{p}e^{-j2\pi f_{c}\tau_{p}}$, $y_{pb}^{I}(t)$ denotes other bandpass signals or interference, and $z_{pb}(t)$ is the bandpass noise. To obtain $y(t)$, $y_{pb}(t)$ must first be down-converted and then passed through a low-pass filter to reject out-of-band noise and signals including $y_{pb}^{I}(t)$ and the high-frequency component around $2f_{c}$. According to the established communication theory[13, 14, 15], the noise $z(t)$ is modeled as having infinite bandwidth and therefore infinite power. In other words, $y(t)$ in (41) is an idealized model that has not yet accounted for the effects of low-pass or receive filtering.
Substituting $x_{\rm AF}^{(a)}(t)$ in (36) as $x(t)$ into (41) and comparing with (42), one can see that the time-varying channel tap in (43) is obtained by directly sampling $y(t)$ with an interval of $\frac{T}{N}$, where the pulse-shaping filter $a(t)$ is a unit-amplitude Nyquist pulse. Therefore, the effect of practically required low-pass or receive filtering is not considered in (42), and the statistical characteristics of of $\mathbf{z}$ may not be well-justified.
Recall that the approximate aliased chirps in (38) are orthogonal to each other when $a(t)$ is a root-Nyquist pulse-shaping filter. Maintaining this orthogonality at the output of the receive filter under an ideal channel naturally suggests employing the same root Nyquist pulse-shaping filter as the receive filter, which also serves as the matched filter.
Let us apply the matched filter $a^{*}(-t)$ to $y(t)$. The sampled output of the matched filter is given by
$y^{\rm MF}[k^{\prime}]$ $=\sum_{l}h_{k^{\prime},l}^{\rm MF}x_{\rm AF}[k^{\prime}-l]+z^{\rm MF}[k^{\prime}],$ (49)
where (see Appendix B for a detailed derivation)
$h_{k^{\prime},l}^{\rm MF}=\sum_{p=1}^{P}h_{p}\,e^{j2\pi\nu_{p}(k^{\prime}-l)\frac{T}{N}}\,\mathcal{A}_{a}(l\frac{T}{N}+\tau_{1}-\tau_{p},\nu_{p}),$ (50)
which reduces to the classical equivalent channel tap for an LTI channel in [2] when $\nu_{p}=0$ for all $p$.
We can now see that, in a practical receiver, the time-varying channel tap in (50) does not have the elegant form assumed in (43), even with on-grid delays, as $\mathcal{A}_{a}(l\frac{T}{N},\nu_{p})$ is generally not zero for $\nu_{p}\neq 0$. In other words, the effective channel cannot be represented by a pure DD spreading function. As a result, the ideal I/O relation in (47) does not exist in practice, thereby rendering the subsequent discussion and conclusions based on it less relevant to practical systems.
Similarly, after discarding the CPP, (49) can be written in matrix form as
$\mathbf{y}^{\rm MF}$ $=\mathbf{H}^{\rm MF}\mathbf{x}+\mathbf{z}^{\rm MF},$ (51)
where $\mathbf{z}^{\rm MF}\sim\mathcal{CN}(0,N_{0}\mathbf{I})$ since $a(t)$ is a root-Nyquist pulse for the Nyquist interval $\frac{T}{N}$ with unit energy $\int|a(t)|^{2}dt=1$. Further applying the DAFT matrix $\mathbf{A}$ to $\mathbf{y}^{\rm MF}$, we have
$\mathbf{y}_{u}^{\rm MF}$ $=\mathbf{A}\mathbf{y}^{\rm MF}=\underbrace{\mathbf{A}\mathbf{H}^{\rm MF}\mathbf{A}^{\mathsf{H}}}_{\mathbf{H}_{u}^{\rm MF}}\mathbf{x}_{u}+\mathbf{A}\mathbf{z}^{\rm MF},$ (52)
where $\mathbf{H}_{u}^{\rm MF}$ is derived in Appendix C. It is noteworthy that the combination of matched filtering, sampling, and DAFT is equivalent to $N$ correlators based on $\phi_{n}^{(a)}(t)$ in (37), which coincides with (2).
(52) shows the explicit practical I/O relation in the DAFT domain. In particular, $\mathbf{H}_{u}^{\rm MF}$ represents the effective coupling coefficient from the transmitted DAFT-domain symbol $X[n]$ to the observed output symbol $Y[n^{\prime}]$ after pulse shaping, physical channel propagation, matched filtering, sampling, and the folding of the CPP portion. This relation provides a foundation for pilot design and channel equalization for practical AFDM waveforms.
VI Simulation Results
In this section, numerical simulations are provided to verify the theoretical analysis presented in the previous sections. In particular, we validate the derived PSD and the exact I/O relation over DD channels.
In the simulations, the AFDM parameters are set to $T=266.667\mu s$, $N=1024$, and the corresponding subcarrier spacing is $\frac{1}{T}=3.75$ kHz. The transmitted symbols are drawn from a $16$-QAM constellation with average power $\sigma_{X}^{2}=1$. To generate the approximate aliased AFDM waveform, the AFDM sequence is pulse-shaped by an square root raised cosine (SRRC) filter with roll-off factor $\beta$ and span $Q\frac{T}{N}$. The multipath channel follows the standard EVA profile [23], consisting of $P=9$ paths with delays that are off-grid with respect to $\frac{T}{N}$. The carrier frequency is $f_{c}=5$GHz, and the Doppler frequency of the $p$-th path is generated using Jakes’ formula $\nu_{p}=\nu_{\textrm{max}}\cos(\theta_{p})$, where the maximum Doppler $\nu_{\textrm{max}}$ is determined by the terminal speed and $\theta_{p}$ is uniformly distributed over $\pi,\pi]$.
VI-A Verification of PSD
We first verify the derived PSD of ideal AFDM waveform under the developed PS-OFDM framework. Fig. 1 compares the simulated PSD with the theoretical PSD derived in Section III, where the chirp parameters are $c_{1}=\frac{1}{4N}$, $c_{2}=\frac{1}{3N}$, respectively. As shown in the figure, the simulated results closely match the analytical PSD. In particular, the AFDM waveform has a bandwidth of approximately $\frac{2c_{1}N^{2}+N-1}{T}\approx 5.76$MHz, confirming the correctness of the theoretical characterization of the spectral behavior of the ideal AFDM waveform.
Figure 1: PSD of ideal AFDM waveform.
VI-B Verification of I/O relation
Next, we verify the derived I/O relation for implemented AFDM waveforms over DD channels, where the chirp parameters are $c_{1}=1$, $c_{2}=2$, respectively. The accuracy of the theoretical model is evaluated in terms of the normalized mean square error (NMSE) between the received DAFT domain signal $\mathbf{y}^{\rm MF}_{u}$ predicted by the derived I/O relation and the received signal obtained from direct waveform simulation through the continuous-time DD channel.
VI-B1 Impact of Terminal Speed
Fig. 2 shows the NMSE versus terminal speed, which determines the maximum Doppler shift in the DD channel. The terminal speed is varied from 0 to $500$ km/h, while the roll-off factor and filter span are fixed at $\beta=0.2$ and $Q=12$, respectively. The results demonstrate that the derived I/O relation accurately predicts the received signal over a wide range of Doppler conditions. In particular, the NMSE remains consistently below 35-35dB even at high mobility. This indicates that variations in terminal speed do not significantly affect the consistency between the theoretical model and the waveform simulation, confirming that the derived I/O relation maintains a high level of accuracy under highly time-varying channel conditions.
Figure 2: NMSE verus Speed, $\beta=0.2$, $Q=12$.
VI-B2 Impact of Roll-Off Factor
Fig. 3 presents the NMSE performance under different roll-off factors of the sample-wise pulse-shaping filter used in the implemented waveform. The roll-off factor is varied from $0.1$ to $0.9$, while the terminal speed and filter span are fixed at $500$ km/h and $Q=12$, respectively. The results show that the derived I/O relation remains valid across various roll-off factors. In particular, the NMSE consistently remains within the relatively low range of 37-37 dB to 33-33 dB as the roll-off factor varies. Although the roll-off factor has some influence on the error level, its overall impact is limited. These results indicate that the derived I/O relation can still accurately characterize the I/O behavior of the implemented AFDM waveform under different pulse-shaping settings.
Figure 3: NMSE verus $\beta$, 500500km/h, $Q=12$.
VI-B3 Impact of Filter Span
Fig. 4 illustrates the NMSE as a function of the filter span, where $Q$ is varied from $6$ to $20$, while the terminal speed and roll-off factor are fixed at $500$ km/h and $\beta=0.2$, respectively. The results show that the NMSE remains within the range of 36-36 dB to 32-32 dB while exhibiting a clear decreasing trend as the filter span increases. This indicates that a larger filter span provides a more accurate approximation to the ideal pulse-shaping filter $a(t)$, thereby reducing the approximation error and improving the agreement between the theoretical model and the direct waveform simulation results. Moreover, when the span becomes sufficiently large, the improvement in NMSE gradually diminishes, suggesting that the error reduction begins to saturate. These results further confirm that the derived I/O relation accurately captures the behavior of implemented AFDM waveforms.
Figure 4: NMSE verus $Q$, 500500km/h, $\beta=0.2$.
VI-C Implementation Complexity Comparison
We compare the implementation complexity of AFDM with that of ODDM. Note that the implemented AFDM signal has a time span of approximately $T$ and a bandwidth of approximately $\frac{N}{T}$. Then, the parameters of ODDM are configured as follows:
  • Number of subcarriers: $N_{\rm OD}$
  • Number of ODDM symbols: $M_{\rm OD}=\frac{N}{N_{\rm OD}}$
  • Subcarrier spacing: $\frac{1}{T}$
to occupy nearly the same TF region.
The complexity is evaluated in terms of the transform operations and filtering procedures required for waveform generation only. It is known that the ODDM waveform can also be approximately realized by generating $N$ time-domain samples and subsequently passing them through a root-Nyquist pulse-shaping filter [5, 9, 8]. Therefore, the comparison focuses on the generation of the $N$ time-domain samples.
As shown above, AFDM is an $N$-subcarrier system, and the complexity associated with the inverse fast Fourier transform (IFFT) operation is $\mathcal{O}(N\log N)$. In contrast, ODDM is an $N_{\rm OD}$-subcarrier system and requires $M_{\rm OD}$ IFFTs of size $N_{\rm OD}$, resulting in a complexity of
$M_{\rm OD}\mathcal{O}(N_{\rm OD}\log N_{\rm OD})=\mathcal{O}(N\log N_{\rm OD})\ll\mathcal{O}(N\log N),$
because usually $N_{\rm OD}\ll N$. Furthermore, it is known that a CP is not necessary for ODDM signals [12], as ODDM is essentially a pseudo-impulse-based transmission scheme for DD channels [9]. As a result, the implementation complexity of ODDM is significantly lower than that of AFDM.
VII Conclusion
This paper develops a unified framework for chirp-domain waveforms, including OCDM and AFDM. By examining their continuous-time representations, it has been shown that these waveforms naturally fall within the conventional WH framework for multicarrier waveforms, where the root chirp corresponds to the prototype pulse. This perspective further reveals that chirp-domain waveforms can be interpreted as PS-OFDM with a constant-envelope prototype pulse, under which their PSD can be characterized analytically.
Practical implementations based on sub-Nyquist time-domain samples have also been analyzed. The resulting aliased waveforms and the orthogonality properties of the embedded aliased chirps have been investigated, showing that orthogonality can still be preserved in implemented waveforms through appropriate sample-wise pulse shaping. An exact I/O relation over DD channels has been derived, revealing that the effective channel observed at practical receivers does not generally follow the conventional DD spreading-function model. Simulations are provided to verify the derived theoretical results. Future work may extend the proposed framework to receiver design and waveform optimization under off-grid DD channels, as well as to multi-antenna systems.
Appendix A Orthogonality Analysis of Aliased Chirps
Recall that the aliased chirp is given by
$\hat{\phi}_{n}(t)=e^{j2\pi\left(c_{2}n^{2}+\frac{c_{1}}{\Delta t^{2}}t^{2}+\frac{n}{T}t-\frac{q_{n}(t)}{\Delta t}t\right)}$ (53)
where $T=N\Delta t$, $C=2N|c_{1}|$, and the interval index $q_{n}(t)$ updates at boundaries $t_{n,q}=\frac{T}{C}q-\frac{n}{C}\Delta t$. This boundary condition allows us to express $q_{n}(t)$ explicitly using the floor function:
$q_{n}(t)=\left\lfloor\frac{C}{T}t+\frac{n}{N}\right\rfloor.$ (54)
Then, the inner product
$I_{n,n^{\prime}}$ $=\int_{0}^{T}\hat{\phi}_{n}(t)\hat{\phi}_{n^{\prime}}^{*}(t)dt$ (55)
$=e^{j2\pi c_{2}(n^{2}-n^{\prime 2})}\int_{0}^{T}e^{j2\pi\left(\frac{n^{\prime\prime}}{T}t-\frac{\Delta q(t)}{\Delta t}t\right)}dt$ (56)
where $n^{\prime\prime}=n-n^{\prime}$ and $\Delta q(t)=q_{n}(t)-q_{n^{\prime}}(t)$.
We then partition the integration interval $[0,T]$ into $C$ segments of length $T/C$ and introduce a local time variable $\varepsilon\in[0,1)$ and an interval index $c\in\{0,1,\dots,C-1\}$, such that $t=(\varepsilon+c)\frac{T}{C}$.
Substitute this into $\Delta q(t)$. Because $c$ is an integer, we can use the floor function property $\lfloor x+c\rfloor=\lfloor x\rfloor+c$ to factor $c$ out:
$\Delta q(t)$ $=\left\lfloor(\varepsilon+c)+\frac{n}{N}\right\rfloor-\left\lfloor(\varepsilon+c)+\frac{n^{\prime}}{N}\right\rfloor$
$=\left(\left\lfloor\varepsilon+\frac{n}{N}\right\rfloor+c\right)-\left(\left\lfloor\varepsilon+\frac{n^{\prime}}{N}\right\rfloor+c\right)$
$=\left\lfloor\varepsilon+\frac{n}{N}\right\rfloor-\left\lfloor\varepsilon+\frac{n^{\prime}}{N}\right\rfloor\equiv\Delta q(\varepsilon).$ (57)
The integer $c$ completely cancels out of the expression. Because the resulting expression depends solely on the local time variable $\varepsilon$ and is independent of $c$, we introduce the shorthand $\equiv\Delta q(\varepsilon)$. This explicitly denotes that $\Delta q$ is periodic across every segment $c$.
Substituting $t=(\varepsilon+c)\frac{T}{C}$ into the integral and rewriting as a sum over $c$, we obtain
$I_{n,n^{\prime}}\propto\frac{T}{C}\sum_{c=0}^{C-1}\int_{0}^{1}e^{j2\pi\left(\frac{n^{\prime\prime}}{T}-\frac{\Delta q(\varepsilon)}{\Delta t}\right)(\varepsilon+c)\frac{T}{C}}d\varepsilon.$ (58)
Using $T=N\Delta t$, the phase simplifies to $\frac{n^{\prime\prime}-N\Delta q(\varepsilon)}{C}(\varepsilon+c)$. We define the integer function $\eta(\varepsilon)=n^{\prime\prime}-N\Delta q(\varepsilon)$, allowing us to separate the variables:
$I_{n,n^{\prime}}\propto\frac{T}{C}\int_{0}^{1}e^{j2\pi\frac{\eta(\varepsilon)}{C}\varepsilon}\underbrace{\left[\sum_{c=0}^{C-1}e^{j2\pi\frac{\eta(\varepsilon)}{C}c}\right]}_{\text{Geometric Sum }S(\varepsilon)}d\varepsilon,$ (59)
where $\eta(\varepsilon)=(n-n^{\prime})-N\left(\left\lfloor\varepsilon+\frac{n}{N}\right\rfloor-\left\lfloor\varepsilon+\frac{n^{\prime}}{N}\right\rfloor\right)$ is an integer with $|\eta(\varepsilon)|\leq N-1$.
One can see that the orthogonality depends on the geometric sum $S(\varepsilon)$ in (59). The sum evaluates to $C$ if $\eta(\varepsilon)$ is a multiple of $C$, and 0 otherwise.
Case 1: $C$ is an integer, and $C\geq N$
If $C\geq N$, then $|\eta(\varepsilon)|<C$. Since $n\neq n^{\prime}$, $\eta(\varepsilon)$ cannot be 0. Therefore, $\eta(\varepsilon)$ is never a multiple of $C$. Perfect orthogonality holds.
Case 2: $C$ is an integer, and $C<N$
Because $|\eta(\varepsilon)|$ can reach N1N-1, it is now possible for $\eta(\varepsilon)=\pm C$. When this occurs for specific pairs of chirps (where $n-n^{\prime}$ is a multiple of $C$), $e^{j2\pi\frac{\eta(\varepsilon)}{C}}=1$. Orthogonality fails and aliasing occurs between chirps separated by index $C$.
Appendix B Derivation of the Effective Channel Tap $h_{k^{\prime},l}^{\mathrm{MF}}$
The channel tap $h_{k^{\prime},l}^{\mathrm{MF}}$ can be interpreted as the channel observed when a discrete sequence with symbol interval $T_{s}$, pulse-shaped by $a(t)$ with Nyquist interval $T_{s}$, passes through the DD channel and is subsequently processed by a matched filter.
Consider the transmitted waveform
$x(t)=\sum_{k}x[k]a\left(t-kT_{s}\right).$ (60)
Substituting (60) into (41) and ignoring the noise term yields
$y(t)$ $=\sum_{p=1}^{P}h_{p}e^{j2\pi\nu_{p}(t-\tau_{p})}\sum_{k}x[k]a(t-\tau_{p}-kT_{s}),$
$=\sum_{p=1}^{P}\sum_{k}h_{p}x[k]e^{j2\pi\nu_{p}(t-\tau_{p})}a(t-\tau_{p}-kT_{s}).$
The output of the matched filter is given by
$y^{\mathrm{MF}}(t)$ $=\int_{-\infty}^{\infty}y(\lambda)a^{*}(\lambda-t)d\lambda,$
$=\sum_{p=1}^{P}\sum_{k}h_{p}x[k]e^{-j2\pi\nu_{p}\tau_{p}}$
$\quad\times\int_{-\infty}^{\infty}e^{j2\pi\nu_{p}\lambda}a(\lambda-\tau_{p}-kT_{s})a^{*}(\lambda-t)d\lambda.$
Applying the change of variables $\xi=\lambda-\tau_{p}-kT_{s}$ leads to
$y^{\mathrm{MF}}(t)$ $=\sum_{p=1}^{P}\sum_{k}h_{p}x[k]e^{j2\pi\nu_{p}kT_{s}}$
$\quad\times\int_{-\infty}^{\infty}a(\xi)a^{*}(\xi-(t-\tau_{p}-kT_{s}))e^{j2\pi\nu_{p}\xi}d\xi.$
Letting $\tau=t-\tau_{p}-kT_{s}$, we obtain
$y^{\mathrm{MF}}(t)$ $=\sum_{p=1}^{P}\sum_{k}h_{p}x[k]e^{j2\pi\nu_{p}kT_{s}}\mathcal{A}_{a}(t-\tau_{p}-kT_{s},\nu_{p}),$
where $\mathcal{A}_{a}(\tau,\nu)$ is the ambiguity function of $a(t)$.
Next, sampling at $t=\tau_{1}+k^{\prime}T_{s}$ and defining the delay tap index $l\triangleq k^{\prime}-k$, we obtain
$y^{\mathrm{MF}}[k^{\prime}]$ $\triangleq y^{\mathrm{MF}}(\tau_{1}+k^{\prime}T_{s})$
$=\sum_{l}\sum_{p=1}^{P}h_{p}e^{j2\pi\nu_{p}(k^{\prime}-l)T_{s}}$
$\quad\times\mathcal{A}_{a}(lT_{s}+\tau_{1}-\tau_{p},\nu_{p})x[k^{\prime}-l].$
Accordingly, the effective channel tap is given by
$h_{k^{\prime},l}^{\mathrm{MF}}\triangleq\sum_{p=1}^{P}h_{p}e^{j2\pi\nu_{p}(k^{\prime}-l)T_{s}}\mathcal{A}_{a}(lT_{s}+\tau_{1}-\tau_{p},\nu_{p}).$
Substituting $T_{s}=\frac{T}{N}$ completes the derivation.
Appendix C Derivation of Channel Matrix $\mathbf{H}^{\rm MF}_{u}$
Since $A_{a}(\cdot)$ has a finite support, we may limit the delay index $l$ as $l=0,\cdots,L-1$ and rewrite (49) as
$y^{\mathrm{MF}}[k^{\prime}]$ $=\sum_{l=0}^{L-1}h_{k^{\prime},l}^{\mathrm{MF}}x_{\rm AF}[k^{\prime}-l]+z^{\mathrm{MF}}[k^{\prime}].$ (61)
Then, (61) can be expressed in matrix form as
$\hat{\mathbf{H}}^{\mathrm{MF}}$ $=\hat{\mathbf{H}}^{\mathrm{MF}}\mathbf{x}^{\prime}+\mathbf{z}^{\mathrm{MF}},$ (62)
where $\mathbf{y}^{\mathrm{MF}}=[y^{\mathrm{MF}}[0],y^{\mathrm{MF}}[1],\cdots,y^{\mathrm{MF}}[N-1]]^{\mathsf{T}}$, $\mathbf{x}^{\prime}=[x_{\rm AF}[-L+1],\cdots,x_{\rm AF}[-1],x_{\rm AF}[0],\cdots,x_{\rm AF}[N-1]]^{\mathsf{T}}$, $\hat{\mathbf{H}}^{\mathrm{MF}}\in\mathbb{C}^{N\times(N+L-1)}$ is the time-domain channel matrix with entries
$[\hat{\mathbf{H}}^{\mathrm{MF}}]_{k^{\prime},n}$ $=\begin{cases}h_{k^{\prime},k^{\prime}-n}^{\mathrm{MF}},&0\leq k^{\prime}-n\leq L-1,\\ 0,&\text{otherwise},\end{cases}$ (63)
for $k^{\prime}=0,\cdots,N-1$ and $n=-L+1,\cdots,N-1$, and the negative-index entries in $\mathbf{x}^{\prime}$ corresponds to the CPP.
For $k^{\prime}-l<0$, namely $l=k^{\prime}+1,\cdots,L-1$, let us define
$h_{k^{\prime},l}^{\mathrm{MF}\prime}$ $\triangleq h_{k^{\prime},l}^{\mathrm{MF}}e^{-j2\pi c_{1}\left(N^{2}+2N(k^{\prime}-l)\right)}.$ (64)
After CPP folding, (61) becomes
$y^{\mathrm{MF}}[k^{\prime}]=$ $\sum_{l=0}^{k^{\prime}}h_{k^{\prime},l}^{\mathrm{MF}}x_{\rm AF}[k^{\prime}-l]$
$+\sum_{l=k^{\prime}+1}^{L-1}h_{k^{\prime},l}^{\mathrm{MF}\prime}x_{\rm AF}[N+k^{\prime}-l]+z^{\mathrm{MF}}[k^{\prime}].$
Hence, let $\mathbf{x}=[x_{\rm AF}[0],x_{\rm AF}[1],\cdots,x_{\rm AF}[N-1]]^{\mathsf{T}}$, we obtain the CPP-folded square-matrix representation
$\mathbf{y}^{\mathrm{MF}}$ $=\mathbf{H}^{\rm MF}\mathbf{x}+\mathbf{z}^{\mathrm{MF}},$ (65)
where $\mathbf{H}^{\rm MF}\in\mathbb{C}^{N\times N}$ is placed as
$[\mathbf{H}^{\rm MF}]_{k^{\prime},k^{\prime}-l}$ $=h_{k^{\prime},l}^{\mathrm{MF}},\qquad l=0,\cdots,k^{\prime},$
$[\mathbf{H}^{\rm MF}]_{k^{\prime},N+k^{\prime}-l}$ $=h_{k^{\prime},l}^{\mathrm{MF}\prime},\qquad l=k^{\prime}+1,\cdots,L-1.$ (66)
The DAFT-domain effective channel matrix is given by
$\mathbf{H}_{u}^{\mathrm{MF}}$ $=\mathbf{A}\mathbf{H}^{\rm MF}\mathbf{A}^{\mathsf{H}},$ (67)
where the $(q,k)$-th entry of $\mathbf{H}_{u}^{\mathrm{MF}}$ is shown on the top of the next page.
$[\mathbf{H}_{u}^{\mathrm{MF}}]_{q,k}$ $=\sum_{k^{\prime}=0}^{N-1}\sum_{n=0}^{N-1}\phi^{*}[k^{\prime},q]\,[\mathbf{H}^{\mathrm{MF}}]_{k^{\prime},n}\,\phi[n,k]$
$=\sum_{k^{\prime}=0}^{N-1}\phi^{*}[k^{\prime},q]\left(\sum_{l=0}^{k^{\prime}}h_{k^{\prime},l}^{\mathrm{MF}}\phi[k^{\prime}-l,k]+\sum_{l=k^{\prime}+1}^{L-1}h_{k^{\prime},l}^{\mathrm{MF}\prime}\phi[N+k^{\prime}-l,k]\right)$
$=\frac{1}{N}e^{j2\pi c_{2}(k^{2}-q^{2})}\sum_{k^{\prime}=0}^{N-1}\sum_{l=0}^{L-1}h_{k^{\prime},l}^{\mathrm{MF}}e^{j2\pi c_{1}\left((k^{\prime}-l)^{2}-k^{\prime 2}\right)}e^{j2\pi\left(\frac{(k^{\prime}-l)k-k^{\prime}q}{N}\right)}.$ (68)

### 31 Mar 2026 v2

@@
 In this paper, a unified framework for chirp-domain waveforms, including orthogonal chirp division multiplexing (OCDM) and affine frequency division multiplexing (AFDM), is developed.
-Based on their continuous-time representations, we show that these waveforms fall within the conventional Weyl-Heisenberg (WH) framework for multicarrier (MC) waveforms, where the root chirp corresponds directly to the prototype pulse in the WH framework.
+Based on their continuous-time representations, we show that these waveforms fall within the conventional Weyl-Heisenberg (WH) framework for multicarrier (MC) waveforms, where the root chirp corresponds to the prototype pulse in the WH framework.
 Since the chirp is a constant-envelope signal and is transparent to subcarrier orthogonality, these waveforms can be further interpreted as pulse-shaped (PS) orthogonal frequency division multiplexing (OFDM).
@@
 It is shown that the aliased chirps are conditionally orthogonal, whereas the implemented approximate aliased chirps can maintain mutual orthogonality when an appropriate sample-wise pulse-shaping filter is applied.
-We further derive an exact input-output relation for the implemented chirp-domain waveform over a delay-Doppler (DD) channel, showing that the effective channel observed at a practical receiver does not, in general, admit a DD spreading-function model commonly assumed in the literature.
+We further derive an exact input-output (I/O) relation for the implemented chirp-domain waveform over delay-Doppler (DD) channels, showing that the effective channel at a practical receiver does not, in general, conform to a superposition of pure path-wise DD components, resulting in a non-negligible deviation from the I/O relation commonly used in the literature.
 The implementation complexity is also investigated and compared with that of orthogonal delay-Doppler division multiplexing (ODDM), the DD-domain MC waveform defined within the evolved WH framework.
@@
 Such mobility induces a large Doppler spread and shortens the channel coherence time.
-In conventional orthogonal frequency-division multiplexing (OFDM) systems, this often manifests as rapidly increasing inter-carrier interference (ICI): once the normalized Doppler spread (e.g., the maximum Doppler shift divided by the subcarrier spacing) is no longer sufficiently small, subcarriers may effectively lose their frequency-domain orthogonality.
+In conventional orthogonal frequency division multiplexing (OFDM) systems, this often manifests as rapidly increasing inter-carrier interference (ICI).
+Once the normalized Doppler spread (e.g., the maximum Doppler shift divided by the subcarrier spacing) is no longer sufficiently small, subcarriers may effectively lose their frequency-domain orthogonality.
 This behavior is consistent with the fact that OFDM is typically designed under an approximately linear time-invariant (LTI) channel assumption, whereas under high mobility the physical channel is more appropriately modeled as a linear time-varying (LTV) system.
@@
 Since the physical units of delay and Doppler are time and frequency, respectively, the DD domain is naturally a time-frequency (TF) domain.
-As a result, it has been shown that DD-domain waveforms can be interpreted within the Weyl-Heisenberg (WH) framework for multicarrier (MC) waveforms, where ODDM evolves the existing WH framework by introducing a (bi)orthogonal WH-subset-based design [8, 9].
-Recent studies of ODDM waveforms over general physical channels can also be found in [10, 11, 12].
-On the other hand, although orthogonal chirps play a central role in chirp-based waveforms, a systematic framework for understanding the fundamentals of these orthogonal-chirp-based waveforms is still lacking.
-Moreover, existing studies of chirp-domain waveforms are often based on discrete-time sequences generated by the discrete Fresnel transform (DFnT) [6] or the discrete affine Fourier transform (DAFT)[7], whereas the transmitted waveform is ultimately a continuous-time signal that propagates through a general physical channel with continuous-valued delay and Doppler.
-This discrepancy may lead to a mismatch between discrete-sequence-level analyses and the underlying channel behavior characterized by continuous-valued delay and Doppler.
+As a result, it has been shown that DD-domain waveforms can be interpreted within the well-known Weyl-Heisenberg (WH) framework for multicarrier (MC) waveforms, where ODDM evolves the existing WH framework by introducing a (bi)orthogonal WH-subset-based design [8, 9, 10].
+Recent studies of ODDM waveforms over general physical channels can be found in [11, 12, 13].
+On the other hand, although orthogonal chirps play a central role in chirp-based waveforms, a systematic framework for understanding their fundamental properties remains elusive.
+Furthermore, while some studies have addressed the continuous-time representation of chirp-domain waveforms [6, 14, 15], much of the existing literature is based on discrete-time sequences generated by the discrete Fresnel transform (DFnT) [6] or the discrete affine Fourier transform (DAFT) [7].
+Given that the transmitted waveform is ultimately a continuous-time signal propagating through a physical channel characterized by continuous-valued delay and Doppler, this discrepancy can lead to a non-negligible mismatch between discrete-sequence-level analyses and actual channel behavior.
 In this paper, we develop a unified framework for chirp-domain waveforms, including OCDM and AFDM.
-We show that these waveforms also fall within the conventional WH framework for MC waveforms, where the root chirp corresponds directly to the prototype pulse in the WH framework.
+We show that these waveforms also fall within the conventional WH framework for MC waveforms, where the root chirp corresponds to the prototype pulse in the WH framework.
 Under this interpretation, chirp-domain waveforms can be viewed as pulse-shaped (PS) OFDM with a constant-envelope prototype pulse.
@@
 We show that the aliased chirps are conditionally orthogonal, whereas the approximate aliased chirps used in practical implementations can maintain mutual orthogonality when appropriate sample-wise pulse-shaping filters are applied.
-  • For implemented waveforms containing approximate aliased chirps, we derive the exact I/O relation over DD channels.
-Our analysis shows that the effective channel observed at the receiver does not, in general, admit the conventional DD spreading-function representation commonly assumed in the literature.
+  • For the implemented waveform containing approximate aliased chirps, we derive the exact I/O relation over DD channels.
+Our analysis shows that the effective channel at a practical receiver does not, in general, conform to a superposition of pure path-wise DD components commonly assumed in the literature.
 As a result, compared with existing studies that rely solely on discrete-sequence models and may therefore deviate from the behavior observed in practical transceivers, the derived I/O relation provides a more accurate characterization of the practical performance of chirp-domain waveforms.
-  • Simulations are provided to verify the theoretical analysis, validate the derived PSD of chirp-domain waveforms, and confirm the derived I/O relation.
+  • Simulations are provided to verify the theoretical analysis, validate the derived PSD, and confirm the derived I/O relation.
 We also analyze the implementation complexity of chirp-domain waveforms and compare it with that of ODDM.
-When occupying the same TF resources, the implementation complexity of ODDM is shown to be significantly lower than that of chirp-domain waveforms.
+We show that when transmitting the same number of symbols with the same subcarrier spacing, ODDM exhibits notably lower implementation complexity than chirp-domain waveforms.
 The remainder of this paper is organized as follows.
-In Section II, beginning with a general representation of ditigal modulation waveforms, the WH framework for MC waveforms is introduced.
+In Section II, beginning with a general representation of digital modulation waveforms, the WH framework for MC waveforms is introduced.
 Section III presents OCDM and AFDM in their continuous-time forms.
@@
 In Section IV, practical implementations based on sub-Nyquist samples are investigated; the ideal and approximate aliased chirps are presented, and their orthogonality properties are studied.
-Section V derives the I/O relation of implemented chirp-domain waveforms with approximate aliased chirps over DD channels and verifies the effective channel observed at the receiver.
-Simulation validation and complexity comparison are also provided in Section VI.
+Section V derives the I/O relation of the implemented chirp-domain waveform with approximate aliased chirps over DD channels and identifies the effective channel observed at a practical receiver.
+Simulation validation and complexity comparison are provided in Section VI.
 Finally, Section VII concludes the paper.
-Notations: In this paper, uppercase boldface letters are used to represent matrices, and lowercase boldface letters are used for column vectors.
-The superscript $\mathsf{T}$, $\mathsf{H}$ and * denote the transpose, conjugate transpose, and conjugate operators, respectively.
-Also, $\Pi_{T}(t)$ denotes the rectangular function of unit energy and support of $[0,T)$, while $\mathcal{A}_{g,\tilde{g}}(\tau,\nu)$ stands for the cross ambiguity function between two pulses $g(t)$ and $\tilde{g}(t)$, given by
-$\mathcal{A}_{g,{\tilde{g}}}(\tau,\nu)$ $=\langle g(t),{\tilde{g}}(t-\tau)e^{j2\pi\nu(t-\tau)}\rangle,$
+Notations: Uppercase and lowercase boldface letters represent matrices and column vectors, respectively.
+The superscripts $\mathsf{T}$, $\mathsf{H}$, and * indicate transpose, conjugate transpose, and complex conjugate.
+The operator $\star$ denotes convolution.
+The function $\Pi_{T}(t)$ is a unit-energy rectangular pulse with support $[0,T]$, and $\mathcal{A}_{g,\tilde{g}}(\tau,\nu)$ denotes the cross-ambiguity function between $g(t)$ and $\tilde{g}(t)$, given by
+$\mathcal{A}_{g,\tilde{g}}(\tau,\nu)$ $=\langle g(t),\tilde{g}(t-\tau)e^{j2\pi\nu(t-\tau)}\rangle$
 $=\int_{-\infty}^{+\infty}g(t)\,\tilde{g}^{*}(t-\tau)e^{-j2\pi\nu(t-\tau)}\,dt.$
 II MC Waveform Principles
-Given the symbol index $i\in\mathbb{Z}$, a real-valued passband digital modulation waveform with carrier frequency $f_{c}$ and phase $\theta_{c}$ is typically represented as [13, 14, 15, 9]:
-$x_{pb}(t)=\sum_{i}\sqrt{2}A_{i}\cos(2\pi(f_{c}+f_{i}(t))t+\theta_{c}+\theta_{i})\rho_{i}(t),$
+Given the symbol index $i\in\mathbb{Z}$, a real-valued passband digital modulation waveform with carrier frequency $f_{c}$ and phase $\theta_{c}$ is typically represented as [16, 17, 18]:
+$x_{pb}(t)=\sqrt{2}\sum_{i}A_{i}\cos(2\pi(f_{c}+f_{i}(t))t+\theta_{c}+\theta_{i})\rho_{i}(t),$
 where the amplitude $A_{i}$, the phase $\theta_{i}$, the possibly time-varying frequency $f_{i}(t)$, and the finite-duration pulse $\rho_{i}(t)$ can be used for modulation.
@@
 To avoid inter-symbol interference (ISI), the transmit pulses $\{g_{i}(t),i\in\mathbb{Z}\}$ are required to be mutually orthogonal.
-Hence, a digital modulation waveform is defined by a set of orthogonal pulses, also referred to as basis functions.
-For MC systems, the symbol index $i$ is replaced by a time index $m$ and a frequency index $n$, and the MC waveform is given by [8, 9, 16]:
+Hence, a digital modulation waveform can be defined by a set of orthogonal pulses, also referred to as basis functions.
+For MC systems, the symbol index $i$ is replaced by a time index $m$ and a frequency index $n$, and the MC waveform is given by [9, 10, 19, 20]:
 $x(t)=\sum_{m=0}^{M-1}\sum_{n=-\frac{N}{2}}^{\frac{N}{2}-1}X[m,n]\overbrace{e^{j2\pi n\Delta F(t-m\Delta T)}g(t-m\Delta T)}^{g_{m,n}(t)},$ (3)
@@
 Since $e^{j2\pi n\Delta Ft}$ is often called a subcarrier or tone, the transmit pulses $g_{m,n}(t)$ in (3) are truncated (windowed) or pulse-shaped subcarriers, corresponding to a rectangular or non-rectangular $g(t)$, respectively.
-For such an MC waveform characterized by a TF grid $(\Delta T,\Delta F)$ and a prototype pulse $g(t)$, a design framework based on the WH frame theory is well-known[16], as $g_{m,n}(t)$ in (3) is the WH or Gabor function adopted in TF signal analysis [17].
+For such an MC waveform characterized by a TF grid $(\Delta T,\Delta F)$ and a prototype pulse $g(t)$, a design framework based on the WH frame theory is well-known[19, 20], as $g_{m,n}(t)$ in (3) is the WH or Gabor function adopted in TF signal analysis [21].
 A WH set $\{\Delta T,\Delta F,g(t)\}$ is said to be orthogonal given the inner product:
@@
 $=\sum_{n=0}^{N-1}X[n]e^{\frac{j\pi}{4}}e^{-j\frac{\pi}{N}(k-n)^{2}},$ (10)
-for $0\leq k\leq N-1$, where $e^{\frac{j\pi}{4}}e^{-j\frac{\pi}{N}(k-n)^{2}}$ is exactly the $(k,n)$-th entry of the $N\times N$ IDFnT matrix up to a scaling factor.
+for $0\leq k\leq N-1$, where $e^{\frac{j\pi}{4}}e^{-j\frac{\pi}{N}(k-n)^{2}}$ is exactly the $(k,n)$-th entry of the $N\times N$ inverse DFnT (IDFnT) matrix up to a scaling factor.
 III-B AFDM Waveform
@@
 using the inverse AFT, which is also an AFT with parameters (a,b,c,d)(a,-b,-c,d).
-By sampling (11) over both $t$ and $u$ with intervals $\Delta t$ and $\Delta u$, respectively, the discrete AFT (DAFT) and the corresponding inverse DAFT can be defined under the condition [20]
+By sampling (11) over both $t$ and $u$ with intervals $\Delta t$ and $\Delta u$, respectively, the discrete AFT (DAFT) and the corresponding inverse DAFT (IDAFT) can be defined under the condition [24]
 $\Delta t\Delta u=\frac{2\pi b}{N},$ (14)
-where $N$ denots denotes the number of samples over $u$. For the commonly adopted symmetric sampling grid with $\Delta t=\Delta u$, the transform parameter $b$ follows
+where $N$ denotes the number of samples over $u$. For the commonly adopted symmetric sampling grid with $\Delta t=\Delta u$, the transform parameter $b$ follows
 $b=\frac{N\Delta u^{2}}{2\pi}.$ (15)
-In [7], the inverse DAFT is used to generate the discrete-time AFDM sequence as
+In [7], the IDAFT is used to generate the discrete-time AFDM sequence as
 $x_{\rm{AF}}[k]$ $=\frac{1}{\sqrt{N}}\sum_{n=0}^{N-1}X[n]e^{j2\pi\left(c_{2}n^{2}+\frac{nk}{N}+c_{1}k^{2}\right)}$ (16)
@@
 $=\frac{1}{\sqrt{N}}\Pi_{T}(t)e^{j2\pi\left(c_{2}n^{2}+\frac{n}{N\Delta t}t+c_{1}\left(\frac{t}{\Delta t}\right)^{2}\right).}$ (20)
-Bearing in mind that $\Delta t=\frac{T}{N}$, and comparing (20) with (7), it is clear that OCDM is a special case of AFDM with $c_{1}=\frac{1}{2N}.$ Thus, we hereafter use
+Bearing in mind that $\Delta t=\frac{T}{N}$, and comparing (20) with (7), it is clear that OCDM is a special case of AFDM with $c_{1}=-\frac{1}{2N}.$ Thus, we hereafter use
 $x_{\rm{AF}}(t)=\sum_{n=0}^{N-1}X[n]\dot{\phi}_{n}(t)$ (21)
@@
 In other words, $g_{\rm AF}(t)$ can form an orthogonal WH set over the TF grid $\left(T,\frac{1}{T}\right)$.
-Recall that when $g(t)$ forms an orthogonal WH set over the TF grid $\left(\Delta T,\Delta F\right)$, the MC waveform in (3) becomes a PS-OFDM signal.
-As a result, chirp-domain waveforms not only belong to the WH framework but can also be interpreted as PS-OFDM signals with a constant-envelope prototype pulse.
+Recall that when $g(t)$ is non-rectangular and forms an orthogonal WH set over the TF grid $\left(\Delta T,\Delta F\right)$, the MC waveform in (3) becomes a PS-OFDM signal.
+Consequently, chirp-domain waveforms not only fall within the WH framework but can also be interpreted as PS-OFDM signals with a constant-envelope prototype pulse.
 III-D PSD Analysis
@@
 It can be proved that $x(t)$ is wide-sense cyclostationary with period $\Delta T$.
-Given the time-average autocorrelation function of cyclostationary signal is defined as [14] :
+Given the time-average autocorrelation function of cyclostationary signal defined as [17] :
 $\bar{R}_{x}(\tau)$ $=\frac{1}{\Delta T}\int_{0}^{\Delta T}R_{x}(t,\tau)dt,$
@@
 For the ideal AFDM signal in (22), after a centered re-indexing of the subcarrier index to match the unified formulation in (3), and then substituting $g(t)=g_{\rm AF}(t)$, $\Delta T=T$, $\Delta F=1/T$ into (30), we have
-$S_{\rm AF}(f)=\frac{\sigma_{X}^{2}}{NT}\sum_{n=-\frac{N}{2}}^{\frac{N}{2}-1}\left|G_{\rm AF}\left(f-\frac{n}{T}\right)\right|^{2},$ (31)
-where $e^{j2\pi c_{2}n^{2}}$ is a phase factor that does not affect the second-order statistics of $\dot{X}[n]$, and $G(f)$ is the corresponding prototype pulse spectrum given by
+$S_{\rm AF}(f)=\frac{\sigma_{X}^{2}}{NT}\sum_{n=-\frac{N}{2}}^{\frac{N}{2}-1}\left|G_{\rm AF}\left(f-\frac{n+N/2}{T}\right)\right|^{2},$ (31)
+where $e^{j2\pi c_{2}n^{2}}$ in (22) is a phase factor that does not affect the second-order statistics of $\dot{X}[n]$, and $G(f)$ is the corresponding prototype pulse spectrum given by
 $G_{\rm AF}(f)=\int_{0}^{T}e^{j2\pi c_{1}N^{2}\left(\frac{t}{T}\right)^{2}}e^{-j2\pi ft}\,dt.$ (32)
@@
 Remark 1
-A discrete sequence has infinite bandwidth, and therefore must be pulse-shaped into a analog waveform that is band-limited before being carrier-modulated and transmitted over the air.
+A discrete-time sequence has infinite bandwidth and must therefore be pulse-shaped into a band-limited analog waveform before carrier modulation and over-the-air transmission.
+It should be noted that generating a discrete-time sequence followed by pulse shaping is only one possible implementation method.
+The modulation waveform can also be generated directly without explicitly forming the intermediate sequence.
 Remark 2
-In the context of MC waveforms, pulse shaping can be somewhat confusing because an MC symbol usually spans multiple time-domain samples, whereas in single-carrier modulation one baud-rate sample corresponds to one symbol.
-Pulse shaping in MC waveforms typically involves two steps: first, interpolating discrete-time samples using a filter (equivalent to the sample-wise pulse shaping used in the single-carrier case), and then multiplying the result by the symbol-wise prototype pulse (symbol-wise pulse shaping).
-When the prototype pulse is constant-envelope, such as the rectangular pulse, sample-wise pulse shaping alone may be sufficient, subject to appropriate parameter settings [8, 9].
+In MC waveforms, pulse shaping is more complex because each symbol spans multiple time-domain samples usually generated by the inverse fast Fourier transform (IFFT), unlike single-carrier modulation where each sample corresponds to one symbol.
+It typically involves two steps: sample-wise pulse shaping via interpolation filtering (equivalent to that used in the single-carrier case), and subsequent symbol-wise pulse shaping via multiplication with a prototype pulse.
+When the prototype pulse has a constant envelope (e.g., rectangular), sample-wise pulse shaping alone may suffice with appropriate parameter settings [9, 10].
+Such waveforms are also referred to as filtered OFDM [25].
 After the sample-wise pulse shaping, we have the approximate aliased AFDM waveform
@@
 for $0\leq k\leq N-1$.
-The chirp sequence can be represented as a vector $\bm{\phi}_{n}=[\phi_{n}[0],\cdots,\phi_{n}[N-1]]^{T},$ corresponding to the $n$-th column of the IDAFT matrix.
+The chirp sequence can be represented as a vector $\bm{\phi}_{n}=[\phi_{n}[0],\cdots,\phi_{n}[N-1]]^{\mathsf{T}},$ corresponding to the $n$-th column of the IDAFT matrix.
 Since $\langle\bm{\phi}_{n},\bm{\phi}_{n^{\prime}}\rangle=\delta(n-n^{\prime})$, it is well known that
-$\langle\phi_{n}^{(a)}(t),\phi_{n^{\prime}}^{(a)}(t)\rangle=\delta(n-n^{\prime}),$
-when $a(t)$ is a root-Nyquist pulse for the Nyquist interval $\frac{T}{N}$ [15].
+$\langle\phi_{n}^{(a)}(t),\phi_{n^{\prime}}^{(a)}(t)\rangle=N\delta(n-n^{\prime}),$
+when $a(t)$ is a root-Nyquist pulse with Nyquist interval $\frac{T}{N}$ [18].
 As a result, to maintain orthogonality among these approximated aliased chirps, a root-Nyquist pulse-shaping filter is required.
@@
 For the $p$-th path, let $h_{p}$, $\tau_{p}$, and $\nu_{p}$ denote its complex gain, delay, and Doppler shift, respectively.
-This channel model is considered valid during the channel’s ”stationarity” interval[1, 3, 8].
+This channel model is considered valid during the channel’s “stationary” interval[1, 3, 9].
 Its representation in the DD domain, also known as the DD spreading function, is given by
@@
 $\tau_{1}\leq\tau_{2}\leq\cdots\leq\tau_{P}.$ (40)
-and the delay spread of the channel is given by $\mathcal{T}=\tau_{p}-\tau_{1}<T$. For a transmitted baseband signal $x(t)$, the received baseband signal $r(t)$ is expressed as
+Then, the relative delay and the delay spread of the channel are given by $\tilde{\tau}_{p}=\tau_{p}-\tau_{1}$ and $\tilde{\tau}_{P}<T$, respectively.
+For a transmitted baseband signal $x(t)$, the received baseband signal $y(t)$ is usually expressed as
 $y(t)=\sum_{p=1}^{P}h_{p}e^{j2\pi\nu_{p}(t-\tau_{p})}x(t-\tau_{p})+z(t),$ (41)
@@
 V-A Discrete-Sequence Based I/O Relation
-The I/O relation of AFDM waveform has been derived in [7], where each path is assumed with an on-grid delay that can be normalized to an integer as $l_{p}=\tau_{p}\frac{N}{T}.$ Then, the received samples are given as [7]
+The I/O relation of AFDM waveform has been derived in [7], where each path is assumed with an on-grid delay that can be normalized to an integer as $l_{p}=\tau_{p}\frac{N}{T}.$ Then, the received samples are given as
 $y[k^{\prime}]$ $=\sum_{l=0}^{\infty}h_{k^{\prime},l}\,x_{\rm AF}[k^{\prime}-l]+z[k^{\prime}],$ (42)
 where $z[k^{\prime}]$ is the samples of AWGN and
-$h_{k^{\prime},l}$ $=\sum_{p=1}^{P}h_{p}e^{j2\pi\nu_{p}k^{\prime}}\delta(l-l_{p})$ (43)
+$h_{k^{\prime},l}$ $=\sum_{p=1}^{P}h_{p}e^{j2\pi\nu_{p}k^{\prime}\frac{T}{N}}\delta(l-l_{p})$ (43)
 is the time-varying impulse response of channel at time $k^{\prime}$ and delay $l$.
-A chirp-periodic prefix (CPP) with length $L_{\rm CPP}>\lceil{\mathcal{T}}\frac{N}{T}\rceil-1$ is ususally prepended into $x_{AF}[k]$ in (42).
-After discarding the CPP, (42) can be written in the matrix form
+$x_{AF}[\cdot]$ in (42) usually includes a chirp-periodic prefix (CPP) with length $L_{\rm CPP}>\lceil{\tilde{\tau}_{P}}\frac{N}{T}\rceil-1$.
+In the receiver, after discarding the CPP and letting $\mathbf{y}=[y[0],\cdots,y[N-1]]^{\mathsf{T}}$, $\mathbf{x}=[x_{\rm AF}[0],\cdots,x_{\rm AF}[N-1]]^{\mathsf{T}}$, (42) can be expressed in matrix form as
 $\mathbf{y}$ $=\mathbf{H}\mathbf{x}+\mathbf{z},$ (44)
-with $\mathbf{z}\sim\mathcal{CN}(0,N_{0}\mathbf{I})$,
-$\mathbf{H}=\sum_{p=1}^{P}h_{p}\mathbf{\Gamma}_{p}\mathbf{\Delta}_{\nu_{p}}\mathbf{\Pi}^{l_{p}},$ (45)
-and $\mathbf{\Pi}$ is the $N\times N$ forward cyclic-shift matrix, $\mathbf{\Delta}_{\nu_{p}}\triangleq\mathrm{diag}\left(1,e^{j2\pi\nu_{p}},\cdots,e^{j2\pi(N-1)\nu_{p}}\right)$ and $\mathbf{\Gamma}_{p}$ is a $N\times N$ diagonal matrix given by
+where $\mathbf{z}=[z[0],\cdots,z[N-1]]^{\mathsf{T}}\sim\mathcal{CN}(0,N_{0}\mathbf{I})$, and
+$\mathbf{H}=\sum_{p=1}^{P}h_{p}\mathbf{\Gamma}_{p}\mathbf{\Delta}_{\nu_{p}}\mathbf{\Pi}^{l_{p}}.$ (45)
+In this expression, $\mathbf{\Pi}$ denotes the $N\times N$ forward cyclic-shift matrix, $\mathbf{\Delta}_{\nu_{p}}\triangleq\mathrm{diag}\left(1,e^{j2\pi\nu_{p}},\cdots,e^{j2\pi(N-1)\nu_{p}}\right)$, and $\mathbf{\Gamma}_{p}$ is a $N\times N$ diagonal matrix given by
 $\mathbf{\Gamma}_{p}$ $=\mathrm{diag}\!\left(\begin{cases}e^{-j2\pi c_{1}\left(N^{2}-2N(l_{p}-n)\right)},&n<l_{p}\\ 1,&n\geq l_{p}\end{cases}\right).$ (46)
-By applying the DAFT matrix $\mathbf{A}$ to $\mathbf{y}$, the DAFT domain I/O relation becomes
+By applying the $N\times N$ DAFT matrix $\mathbf{A}$ to $\mathbf{y}$, the DAFT domain I/O relation becomes
 $\mathbf{y}_{u}$ $=\mathbf{A}\mathbf{y}=\underbrace{\mathbf{A}\mathbf{H}\mathbf{A}^{\mathsf{H}}}_{\mathbf{H}_{u}}\mathbf{x}_{u}+\mathbf{A}\mathbf{z}.$ (47)
 where $\mathbf{x}_{u}=[X[0],\cdots,X[N-1]]^{\mathsf{T}}$.
-The DAFT domain I/O matrix $\mathbf{H}_{u}$ plays a fundamental role in performance evaluation and AFDM waveform design, such as in selecting the parameters $c_{1}$ and $c_{2}$ to achieve optimal diversity order and reduced pilot overhead [7].
+The DAFT domain I/O matrix $\mathbf{H}_{u}$ plays a fundamental role in AFDM waveform design and performance evaluation, such as in selecting the parameters $c_{1}$ and $c_{2}$ to achieve optimal diversity order and reduced pilot overhead [7].
 V-B Practical Waveform based I/O Relation
 Although the I/O relation in (47) appears appealing, it is derived from the time-domain relation in (42), which, however, does not match the behavior of a practical receiver processing chain, as will be explained below.
-The passband counterpart of (41) is given by
-$y_{pb}(t)=$ $\sum_{p=1}^{P}\tilde{h}_{p}\cos(2\pi(f_{c}+\nu_{p})(t-\tau_{p}))x(t-\tau_{p})$
-$+y_{pb}^{I}(t)+z_{pb}(t),$ (48)
-where $\tilde{h}_{p}$ is a real-valued gain and $h_{p}=\tilde{h}_{p}e^{-j2\pi f_{c}\tau_{p}}$, $y_{pb}^{I}(t)$ denotes other bandpass signals or interference, and $z_{pb}(t)$ is the bandpass noise.
-To obtain $y(t)$, $y_{pb}(t)$ must first be down-converted and then passed through a low-pass filter to reject out-of-band noise and signals including $y_{pb}^{I}(t)$ and the high-frequency component around $2f_{c}$.
-According to the established communication theory[13, 14, 15], the noise $z(t)$ is modeled as having infinite bandwidth and therefore infinite power.
-In other words, $y(t)$ in (41) is an idealized model that has not yet accounted for the effects of low-pass or receive filtering.
-Substituting $x_{\rm AF}^{(a)}(t)$ in (36) as $x(t)$ into (41) and comparing with (42), one can see that the time-varying channel tap in (43) is obtained by directly sampling $y(t)$ with an interval of $\frac{T}{N}$, where the pulse-shaping filter $a(t)$ is a unit-amplitude Nyquist pulse.
-Therefore, the effect of practically required low-pass or receive filtering is not considered in (42), and the statistical characteristics of of $\mathbf{z}$ may not be well-justified.
+In practice, the received real-valued passband signal is
+$y_{pb}(t)=y_{pb}^{S}(t)+y_{pb}^{I}(t)+z_{pb}(t),$ (48)
+where the desired signal component is
+$y_{pb}^{S}(t)$ $=\sqrt{2}\sum_{p=1}^{P}\tilde{h}_{p}\cos\!\left(2\pi(f_{c}+\nu_{p})(t-\tau_{p})\right)x(t-\tau_{p})$
+$=\Re\left\{y^{S}(t)\sqrt{2}e^{j2\pi f_{c}t}\right\},$ (49)
+and $y^{S}(t)=\sum_{p=1}^{P}h_{p}e^{j2\pi\nu_{p}(t-\tau_{p})}x(t-\tau_{p})$ is its complex baseband equivalent with respect to the system carrier reference $\sqrt{2}e^{j2\pi f_{c}t}$.
+Here, $\tilde{h}_{p}$ is a real-valued path gain, $h_{p}=\tilde{h}_{p}e^{-j2\pi f_{c}\tau_{p}}$ is the corresponding complex baseband gain, $y_{pb}^{I}(t)$ denotes the out-of-band interference component, and $z_{pb}(t)$ is passband AWGN with a two-sided PSD of $\frac{N_{0}}{2}$.
+To obtain the baseband signal, $y_{pb}(t)$ is first filtered by a real-valued bandpass filter (BPF) $r_{pb}(t)$ to remove $y^{I}_{pb}(t)$ and is then down-converted[26].
+Let $r(t)$ denote the baseband equivalent of $r_{pb}(t)$.
+After mixing, we obtain
+$\tilde{y}(t)=$ $y^{S}(t)\star r(t)+(y^{S}(t)\star r(t))^{*}e^{-j4\pi f_{c}t}$
+$+z(t)\star r(t)+(z(t)\star r(t))^{*}e^{-j4\pi f_{c}t},$ (50)
+where $z(t)$ is the baseband equivalent of $z_{pb}(t)$. Further applying a low-pass filter (LPF) $\tilde{r}(t)$ to remove the components around $2f_{c}$ yields
+$y(t)$ $=y^{S}(t)\star r(t)\star\tilde{r}(t)+z(t)\star r(t)\star\tilde{r}(t),$ (51)
+where the joint effect of the BPF and the LPF, given by $r(t)\star\tilde{r}(t)$, can be interpreted as an equivalent baseband receive filter[9, 27].
+In classical communication theory [16, 17, 18], the baseband noise $z(t)$ is often modeled as having infinite bandwidth and, consequently, infinite power for mathematical convenience.
+To justify this model, a subsequent filtering stage is typically introduced, which aligns exactly with (51).
+Comparing (41) with (51), it is clear that (41) represents an idealized model that neglects the practical receive filtering mentioned above.
+Substituting (36) into (41), it can be seen that the time-varying channel tap $h_{k^{\prime},l}$ in (42) is a superposition of pure path-wise DD components, obtained by sampling $y(t)$ with interval $\frac{T}{N}$ for a unit-amplitude Nyquist pulse $a(t)$.
+Therefore, the effect of inevitable receive filtering is not captured in (42), and the statistical characterization of $z[k^{\prime}]$ may not be well justified.
 Recall that the approximate aliased chirps in (38) are orthogonal to each other when $a(t)$ is a root-Nyquist pulse-shaping filter.
-Maintaining this orthogonality at the output of the receive filter under an ideal channel naturally suggests employing the same root Nyquist pulse-shaping filter as the receive filter, which also serves as the matched filter.
-Let us apply the matched filter $a^{*}(-t)$ to $y(t)$. The sampled output of the matched filter is given by
+Maintaining this orthogonality at the output of the receive filter under an ideal channel naturally suggests employing the same root-Nyquist pulse-shaping filter as the receive filter, which also serves as the matched filter and the anti-aliasing filter.
+Let us apply the matched filter $a^{*}(-t)$ to $y(t)$, and let $\mathcal{A}_{a,a}(\cdot)$ denote the ambiguity function of $a(t)$. The sampled output of the matched filter is given by
 $y^{\rm MF}[k^{\prime}]$ $=\sum_{l}h_{k^{\prime},l}^{\rm MF}x_{\rm AF}[k^{\prime}-l]+z^{\rm MF}[k^{\prime}],$ (52)
 where (see Appendix B for a detailed derivation)
-$h_{k^{\prime},l}^{\rm MF}=\sum_{p=1}^{P}h_{p}\,e^{j2\pi\nu_{p}(k^{\prime}-l)\frac{T}{N}}\,\mathcal{A}_{a}(l\frac{T}{N}+\tau_{1}-\tau_{p},\nu_{p}),$ (50)
+$\hskip-8.53581pth_{k^{\prime},l}^{\rm MF}=\sum_{p=1}^{P}h_{p}\,e^{j2\pi\nu_{p}(k^{\prime}\frac{T}{N}-\tilde{\tau}_{p})}\,\mathcal{A}_{a,a}(l\frac{T}{N}-\tilde{\tau}_{p},-\nu_{p}),$ (53)
 which reduces to the classical equivalent channel tap for an LTI channel in [2] when $\nu_{p}=0$ for all $p$.
-We can now see that, in a practical receiver, the time-varying channel tap in (50) does not have the elegant form assumed in (43), even with on-grid delays, as $\mathcal{A}_{a}(l\frac{T}{N},\nu_{p})$ is generally not zero for $\nu_{p}\neq 0$.
-In other words, the effective channel cannot be represented by a pure DD spreading function.
-As a result, the ideal I/O relation in (47) does not exist in practice, thereby rendering the subsequent discussion and conclusions based on it less relevant to practical systems.
-Similarly, after discarding the CPP, (49) can be written in matrix form as
+Note that for other pairs of transmit and receive pulses/filters, the equivalent channel taps can be obtained by replacing $\mathcal{A}_{a,a}(\cdot)$ in (53) with the appropriate cross-ambiguity function.
+We can now see that, in a practical receiver, the time-varying channel tap in (53) generally does not take the simplified form assumed in (43), even with on-grid delays, because the term $\mathcal{A}_{a,a}(l\frac{T}{N},-\nu_{p})$ is typically non-zero for $\nu_{p}\neq 0$.
+Consequently, the effective channel generally cannot be expressed as a superposition of pure path-wise DD components.
+This implies that the I/O relation in (47) generally does not hold in practice, which may limit the relevance of discussions and conclusions based solely on that model.
+Remark 3
+For LTI channels with $\nu_{p}=0$ for all $p$, the channel model in (43) remains valid without explicitly accounting for the receive filter, as its effect can be fully absorbed into the equivalent on-grid channel.
+Provided that the equivalent channel remains LTI, which is typically unknown and must be estimated using pilot signals in practice, all analyses and conclusions based on the on-grid LTI channel assumption continue to hold.
+Similarly, after discarding the CPP and letting $\mathbf{y}^{\rm MF}=[y^{\rm MF}[0],\cdots,y^{\rm MF}[N-1]]^{\mathsf{T}}$, (52) can be written in matrix form as
 $\mathbf{y}^{\rm MF}$ $=\mathbf{H}^{\rm MF}\mathbf{x}+\mathbf{z}^{\rm MF},$ (54)
-where $\mathbf{z}^{\rm MF}\sim\mathcal{CN}(0,N_{0}\mathbf{I})$ since $a(t)$ is a root-Nyquist pulse for the Nyquist interval $\frac{T}{N}$ with unit energy $\int|a(t)|^{2}dt=1$.
+where $\mathbf{z}^{\rm MF}=[z^{\rm MF}[0],\cdots,z^{\rm MF}[N-1]]^{\mathsf{T}}\sim\mathcal{CN}(0,N_{0}\mathbf{I})$ because $a(t)$ is a root-Nyquist pulse for the Nyquist interval $\frac{T}{N}$ and has unit energy $\int|a(t)|^{2}dt=1$.
 Further applying the DAFT matrix $\mathbf{A}$ to $\mathbf{y}^{\rm MF}$, we have
@@
 It is noteworthy that the combination of matched filtering, sampling, and DAFT is equivalent to $N$ correlators based on $\phi_{n}^{(a)}(t)$ in (37), which coincides with (2).
-(52) shows the explicit practical I/O relation in the DAFT domain.
-In particular, $\mathbf{H}_{u}^{\rm MF}$ represents the effective coupling coefficient from the transmitted DAFT-domain symbol $X[n]$ to the observed output symbol $Y[n^{\prime}]$ after pulse shaping, physical channel propagation, matched filtering, sampling, and the folding of the CPP portion.
-This relation provides a foundation for pilot design and channel equalization for practical AFDM waveforms.
+(55) provides the explicit practical I/O relation in the DAFT domain.
+Specifically, the effective coupling coefficients between $\mathbf{x}_{u}$ and $\mathbf{y}_{u}^{\rm MF}$, represented by $\mathbf{H}_{u}^{\rm MF}$, account for the entire signal chain, including the IDAFT transformation, pulse shaping, physical channel propagation, receive filtering, sampling, CPP folding, and the final DAFT transformation.
+Since this relation exhibits a non-negligible deviation from $\mathbf{H}_{u}$ in (47), (55) serves as a necessary foundation for pilot design and channel equalization in practical AFDM systems.
 VI Simulation Results
@@
 In particular, we validate the derived PSD and the exact I/O relation over DD channels.
-In the simulations, the AFDM parameters are set to $T=266.667\mu s$, $N=1024$, and the corresponding subcarrier spacing is $\frac{1}{T}=3.75$ kHz.
-The transmitted symbols are drawn from a $16$-QAM constellation with average power $\sigma_{X}^{2}=1$.
-To generate the approximate aliased AFDM waveform, the AFDM sequence is pulse-shaped by an square root raised cosine (SRRC) filter with roll-off factor $\beta$ and span $Q\frac{T}{N}$.
-The multipath channel follows the standard EVA profile [23], consisting of $P=9$ paths with delays that are off-grid with respect to $\frac{T}{N}$.
-The carrier frequency is $f_{c}=5$GHz, and the Doppler frequency of the $p$-th path is generated using Jakes’ formula $\nu_{p}=\nu_{\textrm{max}}\cos(\theta_{p})$, where the maximum Doppler $\nu_{\textrm{max}}$ is determined by the terminal speed and $\theta_{p}$ is uniformly distributed over $\pi,\pi]$.
+[…]
 VI-A Verification of PSD
@@
 VI-B Verification of I/O relation
-Next, we verify the derived I/O relation for implemented AFDM waveforms over DD channels, where the chirp parameters are $c_{1}=1$, $c_{2}=2$, respectively.
+Next, we verify the derived I/O relation for implemented AFDM waveforms over DD channels, using the same chirp parameters $c_{1}=\frac{1}{4N}$ and $c_{2}=\frac{1}{3N}$.
 The accuracy of the theoretical model is evaluated in terms of the normalized mean square error (NMSE) between the received DAFT domain signal $\mathbf{y}^{\rm MF}_{u}$ predicted by the derived I/O relation and the received signal obtained from direct waveform simulation through the continuous-time DD channel.
@@
 The results demonstrate that the derived I/O relation accurately predicts the received signal over a wide range of Doppler conditions.
-In particular, the NMSE remains consistently below 35-35dB even at high mobility.
+In particular, the NMSE remains consistently below 50-50 dB even at high mobility.
 This indicates that variations in terminal speed do not significantly affect the consistency between the theoretical model and the waveform simulation, confirming that the derived I/O relation maintains a high level of accuracy under highly time-varying channel conditions.
@@
 Fig.
-3 presents the NMSE performance under different roll-off factors of the sample-wise pulse-shaping filter used in the implemented waveform.
-The roll-off factor is varied from $0.1$ to $0.9$, while the terminal speed and filter span are fixed at $500$ km/h and $Q=12$, respectively.
-The results show that the derived I/O relation remains valid across various roll-off factors.
-In particular, the NMSE consistently remains within the relatively low range of 37-37 dB to 33-33 dB as the roll-off factor varies.
-Although the roll-off factor has some influence on the error level, its overall impact is limited.
-These results indicate that the derived I/O relation can still accurately characterize the I/O behavior of the implemented AFDM waveform under different pulse-shaping settings.
+3 shows the NMSE performance across various roll-off factors for the pulse-shaping filter used in the implemented waveform.
+With the terminal speed and filter span fixed at $500$ km/h and $Q=12$, respectively, the NMSE ranges from 39-39 dB to 62-62 dB.
+The results exhibit a clear decreasing trend as the roll-off factor increases, indicating that, for a fixed filter span, the pulse-shaping filter $a(t)$ with a larger roll-off factor can be approximated more accurately.
+This reduces truncation error and improves the agreement between the theoretical model and the waveform simulation.
+Moreover, as the roll-off factor becomes sufficiently large, the improvement in NMSE gradually diminishes, suggesting that the error reduction begins to saturate.
+These results demonstrate that the derived I/O relation accurately characterizes the behavior of AFDM waveforms under different pulse-shaping settings.
 Figure 3: NMSE verus $\beta$, 500500 km/h, $Q=12$.
@@
 Fig.
-4 illustrates the NMSE as a function of the filter span, where $Q$ is varied from $6$ to $20$, while the terminal speed and roll-off factor are fixed at $500$ km/h and $\beta=0.2$, respectively.
-The results show that the NMSE remains within the range of 36-36 dB to 32-32 dB while exhibiting a clear decreasing trend as the filter span increases.
-This indicates that a larger filter span provides a more accurate approximation to the ideal pulse-shaping filter $a(t)$, thereby reducing the approximation error and improving the agreement between the theoretical model and the direct waveform simulation results.
-Moreover, when the span becomes sufficiently large, the improvement in NMSE gradually diminishes, suggesting that the error reduction begins to saturate.
-These results further confirm that the derived I/O relation accurately captures the behavior of implemented AFDM waveforms.
+4 illustrates the NMSE as a function of the filter span $Q$, which is varied from $6$ to $20$, with the terminal speed and roll-off factor fixed at $500$ km/h and $\beta=0.2$, respectively.
+The NMSE decreases from 40-40 dB to 57-57 dB as $Q$ increases, since a larger span provides a more accurate approximation of $a(t)$ for a fixed roll-off factor.
+Similar to the previous observation, for sufficiently large filter span, the improvement in NMSE becomes marginal, indicating diminishing returns in error reduction.
+These results further confirm that the derived I/O relation accurately captures the behavior of implemented AFDM waveforms under different truncation lengths.
 Figure 4: NMSE verus $Q$, 500500 km/h, $\beta=0.2$.
@@
 We compare the implementation complexity of AFDM with that of ODDM.
-Note that the implemented AFDM signal has a time span of approximately $T$ and a bandwidth of approximately $\frac{N}{T}$.
-Then, the parameters of ODDM are configured as follows:
-  • Number of subcarriers: $N_{\rm OD}$
-  • Number of ODDM symbols: $M_{\rm OD}=\frac{N}{N_{\rm OD}}$
-  • Subcarrier spacing: $\frac{1}{T}$
-to occupy nearly the same TF region.
+Note that the implemented AFDM signal occupies a time-bandwidth product of approximately $T\times\frac{N}{T}=N$.
+This provides $N$ degrees of freedom (DoF) to carry $N$ digital symbols, which is consistent with classical communication theory [2, 18, 29].
+This also explains why AFDM is implemented using sub-Nyquist samples: the ideal waveform occupies excessive bandwidth for transmitting only $N$ symbols, leading to spectral efficiency well below the available DoF.
+For a fair comparison, the number of subcarriers $N_{\rm OD}$ and the number of symbols $M_{\rm OD}$ in ODDM are configured as $M_{\rm OD}N_{\rm OD}=N$, so that the two systems transmit $N$ digital symbols with the same subcarrier spacing $\frac{1}{T}$, while occupying nearly the same TF region.
 The complexity is evaluated in terms of the transform operations and filtering procedures required for waveform generation only.
@@
 Therefore, the comparison focuses on the generation of the $N$ time-domain samples.
-As shown above, AFDM is an $N$-subcarrier system, and the complexity associated with the inverse fast Fourier transform (IFFT) operation is $\mathcal{O}(N\log N)$.
-In contrast, ODDM is an $N_{\rm OD}$-subcarrier system and requires $M_{\rm OD}$ IFFTs of size $N_{\rm OD}$, resulting in a complexity of
-$M_{\rm OD}\mathcal{O}(N_{\rm OD}\log N_{\rm OD})=\mathcal{O}(N\log N_{\rm OD})\ll\mathcal{O}(N\log N),$
-because usually $N_{\rm OD}\ll N$.
+As shown above, AFDM is an $N$-subcarrier system with IFFT complexity $\mathcal{O}(N\log N)$.
+By comparison, ODDM is an $N_{\rm OD}$-subcarrier system requiring $M_{\rm OD}$ IFFTs of size $N_{\rm OD}$, resulting in a complexity of $M_{\rm OD}\mathcal{O}(N_{\rm OD}\log N_{\rm OD})=\mathcal{O}(N\log N_{\rm OD}),$ which is substantially lower than $\mathcal{O}(N\log N)$, since typically $N_{\rm OD}\ll N$.
 Furthermore, it is known that a CP is not necessary for ODDM signals [13], as ODDM is essentially a pseudo-impulse-based transmission scheme for DD channels [10].
-As a result, the implementation complexity of ODDM is significantly lower than that of AFDM.
+Consequently, when transmitting the same number of digital symbols with the same subcarrier spacing (frequency resolution), the implementation complexity of ODDM is notably lower than that of AFDM.
 VII Conclusion
 This paper develops a unified framework for chirp-domain waveforms, including OCDM and AFDM.
-By examining their continuous-time representations, it has been shown that these waveforms naturally fall within the conventional WH framework for multicarrier waveforms, where the root chirp corresponds to the prototype pulse.
+By examining their continuous-time representations, it is shown that these waveforms naturally fall within the conventional WH framework for multicarrier waveforms, where the root chirp corresponds to the prototype pulse.
 This perspective further reveals that chirp-domain waveforms can be interpreted as PS-OFDM with a constant-envelope prototype pulse, under which their PSD can be characterized analytically.
-Practical implementations based on sub-Nyquist time-domain samples have also been analyzed.
-The resulting aliased waveforms and the orthogonality properties of the embedded aliased chirps have been investigated, showing that orthogonality can still be preserved in implemented waveforms through appropriate sample-wise pulse shaping.
-An exact I/O relation over DD channels has been derived, revealing that the effective channel observed at practical receivers does not generally follow the conventional DD spreading-function model.
-Simulations are provided to verify the derived theoretical results.
-Future work may extend the proposed framework to receiver design and waveform optimization under off-grid DD channels, as well as to multi-antenna systems.
+Practical implementations based on sub-Nyquist time-domain samples are also analyzed.
+The resulting aliased waveforms and their orthogonality properties show that orthogonality can be preserved through appropriate sample-wise pulse shaping.
+Furthermore, the derived I/O relation over DD channels reveals that the effective channel at a practical receiver generally cannot be represented as a superposition of pure path-wise DD components, leading to a non-negligible deviation from the I/O relation commonly used in the literature.
+Simulations are provided to validate the theoretical results.
+Future work may extend these findings to receiver design and waveform optimization for off-grid DD channels, as well as to multi-antenna systems.
 Appendix A Orthogonality Analysis of Aliased Chirps
 Recall that the aliased chirp is given by
-$\hat{\phi}_{n}(t)=e^{j2\pi\left(c_{2}n^{2}+\frac{c_{1}}{\Delta t^{2}}t^{2}+\frac{n}{T}t-\frac{q_{n}(t)}{\Delta t}t\right)}$ (53)
-where $T=N\Delta t$, $C=2N|c_{1}|$, and the interval index $q_{n}(t)$ updates at boundaries $t_{n,q}=\frac{T}{C}q-\frac{n}{C}\Delta t$.
-This boundary condition allows us to express $q_{n}(t)$ explicitly using the floor function:
+$\hat{\phi}_{n}(t)=e^{j2\pi\left(c_{2}n^{2}+\frac{c_{1}}{\Delta t^{2}}t^{2}+\frac{n}{T}t-\frac{q}{\Delta t}t\right)}$ (56)
+where $T=N\Delta t$, $C=2N|c_{1}|$, and the interval index $q$ updates at boundaries $t_{n,q}=\frac{T}{C}q-\frac{n}{C}\Delta t$.
+This boundary condition allows us to express $q$ as an explicit function of $t$, denoted $q_{n}(t)$, using the floor function
 $q_{n}(t)=\left\lfloor\frac{C}{T}t+\frac{n}{N}\right\rfloor.$ (57)
 Then, the inner product
-$I_{n,n^{\prime}}$ $=\int_{0}^{T}\hat{\phi}_{n}(t)\hat{\phi}_{n^{\prime}}^{*}(t)dt$ (55)
+$I_{n,n^{\prime}}$ $=\int_{0}^{T}\hat{\phi}_{n}(t)\hat{\phi}_{n^{\prime}}^{*}(t)dt$
 $=e^{j2\pi c_{2}(n^{2}-n^{\prime 2})}\int_{0}^{T}e^{j2\pi\left(\frac{n^{\prime\prime}}{T}t-\frac{\Delta q(t)}{\Delta t}t\right)}dt$ (58)
 where $n^{\prime\prime}=n-n^{\prime}$ and $\Delta q(t)=q_{n}(t)-q_{n^{\prime}}(t)$.
-We then partition the integration interval $[0,T]$ into $C$ segments of length $T/C$ and introduce a local time variable $\varepsilon\in[0,1)$ and an interval index $c\in\{0,1,\dots,C-1\}$, such that $t=(\varepsilon+c)\frac{T}{C}$.
-Substitute this into $\Delta q(t)$. Because $c$ is an integer, we can use the floor function property $\lfloor x+c\rfloor=\lfloor x\rfloor+c$ to factor $c$ out:
-$\Delta q(t)$ $=\left\lfloor(\varepsilon+c)+\frac{n}{N}\right\rfloor-\left\lfloor(\varepsilon+c)+\frac{n^{\prime}}{N}\right\rfloor$
-$=\left(\left\lfloor\varepsilon+\frac{n}{N}\right\rfloor+c\right)-\left(\left\lfloor\varepsilon+\frac{n^{\prime}}{N}\right\rfloor+c\right)$
+[…]
+Substitute this into $\Delta q(t)$. Because $\hat{c}$ is an integer, we can use the floor function property $\lfloor x+\hat{c}\rfloor=\lfloor x\rfloor+\hat{c}$ to factor $\hat{c}$ out:
+$\Delta q(t)$ $=\left\lfloor(\varepsilon+\hat{c})+\frac{n}{N}\right\rfloor-\left\lfloor(\varepsilon+\hat{c})+\frac{n^{\prime}}{N}\right\rfloor$
+[…]
 $=\left\lfloor\varepsilon+\frac{n}{N}\right\rfloor-\left\lfloor\varepsilon+\frac{n^{\prime}}{N}\right\rfloor\equiv\Delta q(\varepsilon).$ (59)
-The integer $c$ completely cancels out of the expression.
-Because the resulting expression depends solely on the local time variable $\varepsilon$ and is independent of $c$, we introduce the shorthand $\equiv\Delta q(\varepsilon)$.
-This explicitly denotes that $\Delta q$ is periodic across every segment $c$.
-Substituting $t=(\varepsilon+c)\frac{T}{C}$ into the integral and rewriting as a sum over $c$, we obtain
-$I_{n,n^{\prime}}\propto\frac{T}{C}\sum_{c=0}^{C-1}\int_{0}^{1}e^{j2\pi\left(\frac{n^{\prime\prime}}{T}-\frac{\Delta q(\varepsilon)}{\Delta t}\right)(\varepsilon+c)\frac{T}{C}}d\varepsilon.$ (58)
-Using $T=N\Delta t$, the phase simplifies to $\frac{n^{\prime\prime}-N\Delta q(\varepsilon)}{C}(\varepsilon+c)$.
-We define the integer function $\eta(\varepsilon)=n^{\prime\prime}-N\Delta q(\varepsilon)$, allowing us to separate the variables:
-$I_{n,n^{\prime}}\propto\frac{T}{C}\int_{0}^{1}e^{j2\pi\frac{\eta(\varepsilon)}{C}\varepsilon}\underbrace{\left[\sum_{c=0}^{C-1}e^{j2\pi\frac{\eta(\varepsilon)}{C}c}\right]}_{\text{Geometric Sum }S(\varepsilon)}d\varepsilon,$ (59)
+The integer $\hat{c}$ completely cancels out of the expression.
+Because the resulting expression depends solely on the local time variable $\varepsilon$ and is independent of $\hat{c}$, we introduce the shorthand $\equiv\Delta q(\varepsilon)$.
+This explicitly denotes that $\Delta q$ is periodic across every segment $\hat{c}$.
+Substituting $t=(\varepsilon+\hat{c})\frac{T}{C}$ into the integral and rewriting as a sum over $\hat{c}$, we obtain
+$I_{n,n^{\prime}}\propto\frac{T}{C}\sum_{\hat{c}=0}^{C-1}\int_{0}^{1}e^{j2\pi\left(\frac{n^{\prime\prime}}{T}-\frac{\Delta q(\varepsilon)}{\Delta t}\right)(\varepsilon+\hat{c})\frac{T}{C}}d\varepsilon.$ (60)
+[…]
+$I_{n,n^{\prime}}\propto\frac{T}{C}\int_{0}^{1}e^{j2\pi\frac{\eta(\varepsilon)}{C}\varepsilon}\underbrace{\left[\sum_{\hat{c}=0}^{C-1}e^{j2\pi\frac{\eta(\varepsilon)}{C}\hat{c}}\right]}_{\text{Geometric Sum }S(\varepsilon)}d\varepsilon,$ (61)
 where $\eta(\varepsilon)=(n-n^{\prime})-N\left(\left\lfloor\varepsilon+\frac{n}{N}\right\rfloor-\left\lfloor\varepsilon+\frac{n^{\prime}}{N}\right\rfloor\right)$ is an integer with $|\eta(\varepsilon)|\leq N-1$.
 One can see that the orthogonality depends on the geometric sum $S(\varepsilon)$ in (61). The sum evaluates to $C$ if $\eta(\varepsilon)$ is a multiple of $C$, and 0 otherwise.
-Case 1: $C$ is an integer, and $C\geq N$
+Case 1 ($C\geq N$)
 If $C\geq N$, then $|\eta(\varepsilon)|<C$. Since $n\neq n^{\prime}$, $\eta(\varepsilon)$ cannot be 0. Therefore, $\eta(\varepsilon)$ is never a multiple of $C$. Perfect orthogonality holds.
-Case 2: $C$ is an integer, and $C<N$
+Case 2 ($C<N$)
 Because $|\eta(\varepsilon)|$ can reach N1N-1, it is now possible for $\eta(\varepsilon)=\pm C$.
@@
 Orthogonality fails and aliasing occurs between chirps separated by index $C$.
+Let us define an $N\times N$ matrix $\mathbf{I}$ with entries $I_{n,n^{\prime}}$.
+The simulation results for $|I_{n,n^{\prime}}|$ under $N=32$ are provided below to verify the conditional orthogonality of the aliased chirps.
+Figs.
+5 and 6 show the results for $C=48$ and $C=32$, respectively, corresponding to Case 1 ($C\geq N$).
+As observed, the matrix $\mathbf{I}$ is almost entirely concentrated on the main diagonal, indicating that a significant inner product is obtained only when $n=n^{\prime}$.
+This confirms that distinct aliased chirps remain mutually orthogonal, as predicted for Case 1.
+In contrast, Fig.
+7 displays the results for $C=16$, corresponding to Case 2 ($C<N$).
+Here, $\mathbf{I}$ exhibits two distinct off-diagonal “bright bands” at $|n-n^{\prime}|=16$, exactly matching the chirp-index separation specified by $C$.
+This loss of orthogonality occurs at a structured offset rather than being randomly distributed, aligning with the theoretical predictions for Case 2.
+Figure 5: $|I_{n,n^{\prime}}|$, $C=48$, $N=32$.
+Figure 6: $|I_{n,n^{\prime}}|$, $C=32$, $N=32$.
+Figure 7: $|I_{n,n^{\prime}}|$, $C=16$, $N=32$.
 Appendix B Derivation of the Effective Channel Tap $h_{k^{\prime},l}^{\mathrm{MF}}$
-The channel tap $h_{k^{\prime},l}^{\mathrm{MF}}$ can be interpreted as the channel observed when a discrete sequence with symbol interval $T_{s}$, pulse-shaped by $a(t)$ with Nyquist interval $T_{s}$, passes through the DD channel and is subsequently processed by a matched filter.
-Consider the transmitted waveform
-$x(t)=\sum_{k}x[k]a\left(t-kT_{s}\right).$ (60)
-Substituting (60) into (41) and ignoring the noise term yields
-$y(t)$ $=\sum_{p=1}^{P}h_{p}e^{j2\pi\nu_{p}(t-\tau_{p})}\sum_{k}x[k]a(t-\tau_{p}-kT_{s}),$
-$=\sum_{p=1}^{P}\sum_{k}h_{p}x[k]e^{j2\pi\nu_{p}(t-\tau_{p})}a(t-\tau_{p}-kT_{s}).$
-The output of the matched filter is given by
+The channel tap $h_{k^{\prime},l}^{\mathrm{MF}}$ can be interpreted as the channel observed when a discrete sequence with sample interval $T_{s}$, pulse-shaped by $a(t)$, passes through the DD channel and is subsequently processed by a matched filter.
+Let us consider the transmitted waveform $x(t)=\sum_{k}x[k]a\left(t-kT_{s}\right)$, and substitute it into (41) to yield
+$y(t)$ $=\sum_{p=1}^{P}\sum_{k}h_{p}x[k]e^{j2\pi\nu_{p}(t-\tau_{p})}a(t-\tau_{p}-kT_{s}),$
+where the noise term is ignored. The output of the matched filter is given by
 $y^{\mathrm{MF}}(t)$ $=\int_{-\infty}^{\infty}y(\lambda)a^{*}(\lambda-t)d\lambda,$
@@
 $y^{\mathrm{MF}}(t)$ $=\sum_{p=1}^{P}\sum_{k}h_{p}x[k]e^{j2\pi\nu_{p}kT_{s}}$
-$\quad\times\int_{-\infty}^{\infty}a(\xi)a^{*}(\xi-(t-\tau_{p}-kT_{s}))e^{j2\pi\nu_{p}\xi}d\xi.$
-Letting $\tau=t-\tau_{p}-kT_{s}$, we obtain
-$y^{\mathrm{MF}}(t)$ $=\sum_{p=1}^{P}\sum_{k}h_{p}x[k]e^{j2\pi\nu_{p}kT_{s}}\mathcal{A}_{a}(t-\tau_{p}-kT_{s},\nu_{p}),$
-where $\mathcal{A}_{a}(\tau,\nu)$ is the ambiguity function of $a(t)$.
-Next, sampling at $t=\tau_{1}+k^{\prime}T_{s}$ and defining the delay tap index $l\triangleq k^{\prime}-k$, we obtain
+$\quad\times\int_{-\infty}^{\infty}a(\xi)a^{*}(\xi-(t-\tau_{p}-kT_{s}))e^{j2\pi\nu_{p}\xi}d\xi,$
+$=\sum_{p=1}^{P}\sum_{k}h_{p}x[k]e^{j2\pi\nu_{p}(t-\tau_{p})}\mathcal{A}_{a,a}(\hat{\tau},-\nu),$
+where $\hat{\tau}=t-\tau_{p}-kT_{s}$. Then, sampling at $t=\tau_{1}+k^{\prime}T_{s}$ and defining the delay tap index $l\triangleq k^{\prime}-k$, we obtain
 $y^{\mathrm{MF}}[k^{\prime}]$ $\triangleq y^{\mathrm{MF}}(\tau_{1}+k^{\prime}T_{s})$
-$=\sum_{l}\sum_{p=1}^{P}h_{p}e^{j2\pi\nu_{p}(k^{\prime}-l)T_{s}}$
-$\quad\times\mathcal{A}_{a}(lT_{s}+\tau_{1}-\tau_{p},\nu_{p})x[k^{\prime}-l].$
+$=\sum_{l}\sum_{p=1}^{P}h_{p}e^{j2\pi\nu_{p}(k^{\prime}T_{s}-\tilde{\tau}_{p})}$
+$\quad\times\mathcal{A}_{a,a}(lT_{s}-\tilde{\tau}_{p},-\nu_{p})x[k^{\prime}-l].$
 Accordingly, the effective channel tap is given by
-$h_{k^{\prime},l}^{\mathrm{MF}}\triangleq\sum_{p=1}^{P}h_{p}e^{j2\pi\nu_{p}(k^{\prime}-l)T_{s}}\mathcal{A}_{a}(lT_{s}+\tau_{1}-\tau_{p},\nu_{p}).$
+$h_{k^{\prime},l}^{\mathrm{MF}}\triangleq\sum_{p=1}^{P}h_{p}e^{j2\pi\nu_{p}(k^{\prime}T_{s}-\tilde{\tau}_{p})}\mathcal{A}_{a,a}(lT_{s}-\tilde{\tau}_{p},-\nu_{p}).$
 Substituting $T_{s}=\frac{T}{N}$ completes the derivation.
 Appendix C Derivation of Channel Matrix $\mathbf{H}^{\rm MF}_{u}$
-Since $A_{a}(\cdot)$ has a finite support, we may limit the delay index $l$ as $l=0,\cdots,L-1$ and rewrite (49) as
+Since $A_{a,a}(\cdot)$ has a finite support, we may limit the delay index $l$ as $l=0,\cdots,L-1$ and rewrite (52) as
 $y^{\mathrm{MF}}[k^{\prime}]$ $=\sum_{l=0}^{L-1}h_{k^{\prime},l}^{\mathrm{MF}}x_{\rm AF}[k^{\prime}-l]+z^{\mathrm{MF}}[k^{\prime}].$ (62)
 Then, (62) can be expressed in matrix form as
-$\hat{\mathbf{H}}^{\mathrm{MF}}$ $=\hat{\mathbf{H}}^{\mathrm{MF}}\mathbf{x}^{\prime}+\mathbf{z}^{\mathrm{MF}},$ (62)
+$\mathbf{y}^{\mathrm{MF}}$ $=\hat{\mathbf{H}}^{\mathrm{MF}}\mathbf{x}^{\prime}+\mathbf{z}^{\mathrm{MF}},$ (63)
 where $\mathbf{y}^{\mathrm{MF}}=[y^{\mathrm{MF}}[0],y^{\mathrm{MF}}[1],\cdots,y^{\mathrm{MF}}[N-1]]^{\mathsf{T}}$, $\mathbf{x}^{\prime}=[x_{\rm AF}[-L+1],\cdots,x_{\rm AF}[-1],x_{\rm AF}[0],\cdots,x_{\rm AF}[N-1]]^{\mathsf{T}}$, $\hat{\mathbf{H}}^{\mathrm{MF}}\in\mathbb{C}^{N\times(N+L-1)}$ is the time-domain channel matrix with entries
-$[\hat{\mathbf{H}}^{\mathrm{MF}}]_{k^{\prime},n}$ $=\begin{cases}h_{k^{\prime},k^{\prime}-n}^{\mathrm{MF}},&0\leq k^{\prime}-n\leq L-1,\\ 0,&\text{otherwise},\end{cases}$ (63)
-for $k^{\prime}=0,\cdots,N-1$ and $n=-L+1,\cdots,N-1$, and the negative-index entries in $\mathbf{x}^{\prime}$ corresponds to the CPP.
+$[\hat{\mathbf{H}}^{\mathrm{MF}}]_{k^{\prime},\dot{n}}$ $=\begin{cases}h_{k^{\prime},k^{\prime}-\dot{n}}^{\mathrm{MF}},&0\leq k^{\prime}-\dot{n}\leq L-1,\\ 0,&\text{otherwise},\end{cases}$ (64)
+for $k^{\prime}=0,\cdots,N-1$ and $\dot{n}=-L+1,\cdots,N-1$, and the negative-index entries in $\mathbf{x}^{\prime}$ corresponds to the CPP, which is given by
+$x_{\rm AF}[k]=x_{\rm AF}[N+k]e^{-j2\pi c_{1}(N^{2}+2Nk)}$ (65)
+for $k=-L_{\rm CPP},\dots,-1$ and $L_{\rm CPP}\geq L$.
 For $k^{\prime}-l<0$, namely $l=k^{\prime}+1,\cdots,L-1$, let us define
-$h_{k^{\prime},l}^{\mathrm{MF}\prime}$ $\triangleq h_{k^{\prime},l}^{\mathrm{MF}}e^{-j2\pi c_{1}\left(N^{2}+2N(k^{\prime}-l)\right)}.$ (64)
+$\acute{h}_{k^{\prime},l}^{\mathrm{MF}}$ $\triangleq h_{k^{\prime},l}^{\mathrm{MF}}e^{-j2\pi c_{1}\left(N^{2}+2N(k^{\prime}-l)\right)}.$ (66)
 After CPP folding, (62) becomes
 $y^{\mathrm{MF}}[k^{\prime}]=$ $\sum_{l=0}^{k^{\prime}}h_{k^{\prime},l}^{\mathrm{MF}}x_{\rm AF}[k^{\prime}-l]$
-$+\sum_{l=k^{\prime}+1}^{L-1}h_{k^{\prime},l}^{\mathrm{MF}\prime}x_{\rm AF}[N+k^{\prime}-l]+z^{\mathrm{MF}}[k^{\prime}].$
-Hence, let $\mathbf{x}=[x_{\rm AF}[0],x_{\rm AF}[1],\cdots,x_{\rm AF}[N-1]]^{\mathsf{T}}$, we obtain the CPP-folded square-matrix representation
-$\mathbf{y}^{\mathrm{MF}}$ $=\mathbf{H}^{\rm MF}\mathbf{x}+\mathbf{z}^{\mathrm{MF}},$ (65)
+$+\sum_{l=k^{\prime}+1}^{L-1}\acute{h}_{k^{\prime},l}^{\mathrm{MF}}x_{\rm AF}[N+k^{\prime}-l]+z^{\mathrm{MF}}[k^{\prime}].$
+Hence, let $\mathbf{x}_{\rm AF}=[x_{\rm AF}[0],x_{\rm AF}[1],\cdots,x_{\rm AF}[N-1]]^{\mathsf{T}}$, we obtain the CPP-folded square-matrix representation
+$\mathbf{y}^{\mathrm{MF}}$ $=\mathbf{H}^{\rm MF}\mathbf{x}_{\rm AF}+\mathbf{z}^{\mathrm{MF}},$ (67)
 where $\mathbf{H}^{\rm MF}\in\mathbb{C}^{N\times N}$ is placed as
 $[\mathbf{H}^{\rm MF}]_{k^{\prime},k^{\prime}-l}$ $=h_{k^{\prime},l}^{\mathrm{MF}},\qquad l=0,\cdots,k^{\prime},$
-$[\mathbf{H}^{\rm MF}]_{k^{\prime},N+k^{\prime}-l}$ $=h_{k^{\prime},l}^{\mathrm{MF}\prime},\qquad l=k^{\prime}+1,\cdots,L-1.$ (66)
+$[\mathbf{H}^{\rm MF}]_{k^{\prime},N+k^{\prime}-l}$ $=\acute{h}_{k^{\prime},l}^{\mathrm{MF}},\qquad l=k^{\prime}+1,\cdots,L-1.$ (68)
 The DAFT-domain effective channel matrix is given by
 $\mathbf{H}_{u}^{\mathrm{MF}}$ $=\mathbf{A}\mathbf{H}^{\rm MF}\mathbf{A}^{\mathsf{H}},$ (69)
-where the $(q,k)$-th entry of $\mathbf{H}_{u}^{\mathrm{MF}}$ is shown on the top of the next page.
-$[\mathbf{H}_{u}^{\mathrm{MF}}]_{q,k}$ $=\sum_{k^{\prime}=0}^{N-1}\sum_{n=0}^{N-1}\phi^{*}[k^{\prime},q]\,[\mathbf{H}^{\mathrm{MF}}]_{k^{\prime},n}\,\phi[n,k]$
-$=\sum_{k^{\prime}=0}^{N-1}\phi^{*}[k^{\prime},q]\left(\sum_{l=0}^{k^{\prime}}h_{k^{\prime},l}^{\mathrm{MF}}\phi[k^{\prime}-l,k]+\sum_{l=k^{\prime}+1}^{L-1}h_{k^{\prime},l}^{\mathrm{MF}\prime}\phi[N+k^{\prime}-l,k]\right)$
-$=\frac{1}{N}e^{j2\pi c_{2}(k^{2}-q^{2})}\sum_{k^{\prime}=0}^{N-1}\sum_{l=0}^{L-1}h_{k^{\prime},l}^{\mathrm{MF}}e^{j2\pi c_{1}\left((k^{\prime}-l)^{2}-k^{\prime 2}\right)}e^{j2\pi\left(\frac{(k^{\prime}-l)k-k^{\prime}q}{N}\right)}.$ (68)
+where the $(n^{\prime},n)$-th entry of $\mathbf{H}_{u}^{\mathrm{MF}}$ is shown at the top of the next page.
+$[\mathbf{H}_{u}^{\mathrm{MF}}]_{n^{\prime},n}$ $=\sum_{k^{\prime}=0}^{N-1}\sum_{\dot{n}=0}^{N-1}[\mathbf{A}]_{n^{\prime},k^{\prime}}[\mathbf{H}^{\mathrm{MF}}]_{k^{\prime},\dot{n}}[\mathbf{A}^{\mathsf{H}}]_{\dot{n},n}$
+$=\sum_{k^{\prime}=0}^{N-1}[\mathbf{A}]_{n^{\prime},k^{\prime}}\left(\sum_{l=0}^{k^{\prime}}h_{k^{\prime},l}^{\mathrm{MF}}[\mathbf{A}^{\mathsf{H}}]_{k^{\prime}-l,n}+\sum_{l=k^{\prime}+1}^{L-1}\acute{h}_{k^{\prime},l}^{\mathrm{MF}}[\mathbf{A}^{\mathsf{H}}]_{N+k^{\prime}-l,n}\right)$
+$=\frac{1}{N}e^{j2\pi c_{2}(n^{2}-{n^{\prime}}^{2})}\sum_{k^{\prime}=0}^{N-1}\sum_{l=0}^{L-1}h_{k^{\prime},l}^{\mathrm{MF}}e^{j2\pi c_{1}\left((k^{\prime}-l)^{2}-{k^{\prime}}^{2}\right)}e^{j2\pi\left(\frac{(k^{\prime}-l)n-k^{\prime}n^{\prime}}{N}\right)}$ (70)